### v1

# Improving Assembly Code Performance with Large Language Models via Reinforcement Learning

Abstract
Large language models (LLMs) have demonstrated strong performance across a wide range of programming tasks, yet their potential for code optimization remains underexplored. This work investigates whether LLMs can optimize the performance of assembly code, where fine-grained control over execution enables improvements that are difficult to express in high-level languages. We present a reinforcement learning framework that trains LLMs using Proximal Policy Optimization (PPO), guided by a reward function that considers both functional correctness, validated through test cases, and execution performance relative to the industry-standard compiler gcc -O3. To support this study, we introduce a benchmark of 8,072 real-world programs. Our model, Qwen2.5-Coder-7B-PPO, achieves 96.0% test pass rates and an average speedup of 1.47× over the gcc -O3 baseline, outperforming all 20 other models evaluated, including Claude-3.7-sonnet. These results indicate that reinforcement learning can unlock the potential of LLMs to serve as effective optimizers for assembly code performance.
1 Introduction
Figure 1: Overview of the assembly code optimization task. Given a C program and its baseline assembly from gcc -O3, an LLM is fine-tuned with Proximal Policy Optimization (PPO) to generate improved assembly. The reward function reflects correctness and performance based on test execution.
Recent advances in large language models (LLMs) have achieved state-of-the-art solutions across a wide range of programming tasks [1, 2, 3, 4, 5]. However, their potential for program optimization remains underexplored. Generating highly optimized code is critical in performance-sensitive domains, and prior work has investigated the use of LLMs to optimize C++ and Python programs [6, 7, 8]. In this work, we aim to utilize LLMs to improve the performance of assembly code, extending their capabilities beyond optimization for high-level languages.
Assembly code optimization is traditionally the responsibility of compilers. While modern compilers apply a series of rule-based transformations to improve performance, such a design introduces the classic phase ordering problem [9], where the order of optimizations can substantially affect the performance of the generated code. Due to the inherent complexity of the optimization task, especially the vast space of possible transformation sequences, compilers face fundamental challenges in converging to optimal code, often leaving significant performance on the table [10].
An alternative approach is superoptimization, which searches the space of all programs that are functionally equivalent to the compiler’s output, aiming to identify the most performant variant. In principle, this strategy may yield optimal code. However, the search space grows exponentially with program size, making exhaustive exploration computationally infeasible in practice. Furthermore, prior work on superoptimization [11, 12] has primarily targeted loop-free, straight-line code, where it is more tractable to formally verify that the optimized code is semantically equivalent to the original. As a result, these approaches are not directly applicable to most real-world programs with loops.
In this work, we explore using large language models to optimize the performance of assembly code. Compared with high-level languages such as Python or C++, assembly code operates closer to the hardware, offering fine-grained control over execution and enabling optimizations that are difficult to express or realize in higher-level code. However, this setting poses several challenges. Assembly code is relatively rare and may be underrepresented in pretraining corpora [13], making it harder for LLMs to reason effectively about their behavior. Furthermore, industry compilers such as GCC have been extensively tuned by performance engineers over decades. Achieving additional speedups beyond gcc -O3 (the compiler’s highest optimization level) is a technically challenging task.
To address the challenges of low-level code optimization, we apply reinforcement learning to enhance the ability of LLMs to optimize assembly code. As shown in Figure 1, we use Proximal Policy Optimization (PPO) to train an LLM using a reward function that considers both correctness and performance. Correctness is evaluated based on whether the generated code passes program-specific test cases, and performance is measured by its speedup relative to the baseline produced by gcc -O3. To support this setting, we construct a new dataset of 8,072 assembly programs derived from real-world competitive programming submissions. Each instance includes input-output test cases and baseline assembly code generated by the compiler at the highest optimization level, which serves as the starting point for further optimization.
We evaluate our approach on the proposed benchmark and find that reinforcement learning substantially improves the ability of LLMs to optimize assembly code. Starting from the base model Qwen2.5-Coder-7B-Instruct, which achieves a modest 1.10x speedup over the gcc -O3 baseline, our PPO-trained model reaches 1.47× average speedup and improves both compile and test pass rates to 96.0%. It achieves the strongest performance across all evaluation metrics, outperforming all 20 other models evaluated, including Claude-3.7-sonnet. Ablation studies show that reward functions emphasizing final speedup, rather than intermediate correctness signals, lead to more effective training.
In summary, our contributions are as follows:
  • We introduce the task of optimizing assembly code performance using large language models, aiming for fine-grained performance improvements beyond what traditional compiler optimizations can achieve.
  • We construct a dataset of real-world C programs paired with the corresponding assembly code generated by the gcc -O3 baseline. Using this dataset, we explore improving LLMs on this task through reinforcement learning, applying Proximal Policy Optimization (PPO).
  • We evaluate 21 LLMs on the proposed benchmark and show that our training substantially improves performance: Qwen2.5-Coder-7B-PPO achieves the highest compile and test pass rates, as well as the best average speedup (1.47×) over the gcc -O3 baseline, outperforming all other models (including Claude-3.7-sonnet) across all evaluation metrics.
2 Related Work
Large Language Models for Code.
Benchmarks for evaluating large language models (LLMs) on code generation from natural language specifications have received increasing attention. Notable examples include HumanEval [1], MBPP [2], APPS [3], and more recent efforts [14, 15, 16, 17]. In parallel, many models have been developed to enhance code generation capabilities, such as Codex [1], AlphaCode [18], CodeGen [19], InCoder [20], StarCoder [21], DeepSeek-Coder [22], Code Llama [23], and others [24, 25]. Beyond code generation, LLMs have been applied to real-world software engineering tasks including automated program repair [26, 27], software testing [28, 29], bug localization [30], and transpilation [31, 32]. SWE-bench [4] integrates these tasks into a benchmark for resolving real GitHub issues. Building on this, SWE-agent [5] and subsequent works [33, 34] employ an agent-based framework that leverages LLMs to improve the issue resolution process.
Recent work has also explored LLMs for improving program performance. CodeRosetta [35] targets automatic parallelization, such as translating C++ to CUDA. Other efforts focus on optimizing Python code for efficiency [7, 8] or enabling self-adaptation [36], and improving C++ performance [6]. Of particular relevance are approaches to low-level code optimization [37, 38]. The LLM Compiler foundation models [39, 40] are primarily designed for code size reduction and binary disassembly, whereas our work focuses on optimizing assembly code for performance. LLM-Vectorizer [41] offers a formally verified solution for auto-vectorization, a specific compiler pass. In contrast, our work does not restrict the optimization type and uses test-case validation.
Learning-Based Code Optimization.
The space of code optimization is vast, and many approaches have leveraged machine learning to improve program performance. A classic challenge in compilers is the phase-ordering problem, where performance depends heavily on the sequence of optimization passes. AutoPhase [42] uses deep reinforcement learning to tackle this, while Coreset [43] employs graph neural networks (GNNs) to guide optimization decisions. Modern compilers apply extensive rewrite rules but offer no guarantee of optimality. Superoptimization seeks the most efficient program among all semantically equivalent variants of the compiler output. Traditional methods use stochastic search, such as Markov Chain Monte Carlo [11], with follow-up work improving scalability [44, 12] and extending to broader domains [45, 46]. These rely on formal verification for correctness, restricting them to small, loop-free programs. In contrast, our approach uses test-based validation, enabling optimization of general programs with loops. With the rise of deep learning, substantial attention has turned to optimizing GPU kernel code. AutoTVM [47] pioneered statistical cost model-based search for CUDA code optimization, followed by methods such as Ansor [48], AMOS [49], and other recent systems [50, 51, 52].
More recently, using LLMs as code optimizers has gained popularity [6, 53, 37], with growing attention to reinforcement learning approaches that guide LLMs through reward-based feedback [54, 34]. CodeRL [55] incorporates unit test-based rewards within an actor-critic framework [56], while PPOCoder [57] extends this with Proximal Policy Optimization (PPO) [58], along with other variants [59]. Subsequent efforts have adapted RL-based techniques [60] to additional low-resource programming languages, including Verilog [61]. To the best of our knowledge, our work is the first to apply reinforcement learning to optimize assembly code using LLMs. Assembly code offers fine-grained control and potential for significant performance gains, but it remains underexplored due to limited training data and the complexity of low-level semantics.
3 Methodology
3.1 Task Definition
Let $C$ be a program written in a high-level language such as C. A modern compiler like gcc can compile $C$ into an x86-64 assembly program $P=\texttt{gcc}(C)$, which can then be further assembled into an executable binary. The assembly program $P$ serves as an intermediate representation that exposes low-level optimization opportunities, making it suitable for aggressive performance improvement. We assume the semantics-preserving nature of the compilation process, i.e., $\llbracket C\rrbracket=\llbracket P\rrbracket$, so that the behavior of the assembly program $P$ is identical to that of the source program $C$.
In theory, the goal is to produce a program $P^{\prime}$ that is functionally equivalent to $P$ across the entire input space $\mathcal{X}$, i.e., $P(x)=P^{\prime}(x)$ for all $x\in\mathcal{X}$. Since verifying this property is undecidable in general, we approximate equivalence using a finite test set $\mathcal{T}=\{(x_{i},y_{i})\}_{i=1}^{n}$, where each input-output pair $(x_{i},y_{i})$ captures the expected behavior of $C$.
We say that an assembly program $P^{\prime}$ is valid if it can be successfully assembled and linked into an executable binary. Let $\texttt{valid}(P^{\prime})\in\{\texttt{True},\texttt{False}\}$ denote this property. We define the set of correct programs as:
$\mathcal{S}(P)=\left\{P^{\prime}\mid\texttt{valid}(P^{\prime})\;\land\;\forall(x_{i},y_{i})\in\mathcal{T},\;P^{\prime}(x_{i})=y_{i}\right\}.$
Performance and Speedup.
Let $t(P)$ denote the execution time of $P$ on the test set $\mathcal{T}$, and let $t(P^{\prime})$ be the corresponding execution time for $P^{\prime}$. The speedup of $P^{\prime}$ relative to $P$ is defined as:
$\text{Speedup}(P^{\prime})=\begin{cases}\frac{t(P)}{t(P^{\prime})}&\text{if }P^{\prime}\in\mathcal{S}(P)\text{ and }t(P^{\prime})<t(P),\\ 1&\text{otherwise}.\end{cases}$
Optimization Objective.
The objective is to generate a candidate program $P^{\prime}$ that maximizes $\text{Speedup}(P^{\prime})$. Only programs in $\mathcal{S}(P)$ are eligible for speedup; any candidate that fails to compile into a binary or produces incorrect outputs is assigned a default speedup of 1. This reflects a practical fallback: when the generated program is invalid, the system can revert to the baseline $P$, compiled with gcc -O3, which defines the 1× reference performance. Although $\mathcal{S}(P)$ captures the correctness criteria, we do not restrict the LLM to generate only valid programs. Instead, the model produces arbitrary assembly code, and correctness is verified post hoc via compilation and test execution. We train an LLM using reinforcement learning (see Section 3.3) to generate candidates that both satisfy correctness and achieve performance improvements.
3.2 Dataset Construction
We construct our dataset using C programs from CodeNet [62], a large-scale corpus of competitive programming submissions. Each dataset instance is a tuple $(C,P,\mathcal{T})$, where $C$ is the original C source code, $P=\texttt{gcc\_O3}(C)$ is the corresponding x86-64 assembly generated by compiling $C$ with gcc at the -O3 optimization level, and $\mathcal{T}=\{(x_{i},y_{i})\}_{i=1}^{n}$ is the test set. Since not all CodeNet problems include test inputs, we adopt those provided by prior work [18] to define $x_{i}$, but discard their output labels. Instead, we regenerate each output $y_{i}$ by executing the original submission on input $x_{i}$, as many CodeNet programs are not accepted solutions, and even accepted ones do not reliably pass all test cases.
Given the scale of CodeNet, which contains over 8 million C and C++ submissions, we sample a subset for this study. To focus on performance-critical cases, we sample programs that exhibit the highest relative speedup from gcc -O0 (no optimization) to gcc -O3 (maximum optimization). Such strategy serves two purposes: (1) it favors programs with complex logic that lead to suboptimal performance under -O0 and can be effectively optimized by -O3, and (2) it creates a more challenging setting by starting from code that has already benefited from aggressive compiler optimizations. If an LLM can generate code that further improves upon gcc -O3, it suggests that the model can outperform the compiler’s “expert” solution. The final dataset consists of 7,872 training programs and 200 held-out evaluation programs, with additional statistics provided in Section 4.
3.3 Reinforcement Learning
We conceptualize our task as a standard contextual multi-armed bandit problem [63], defined by a context space $\mathcal{S}$, an action space $\mathcal{A}$, and a reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. Each context $s\in\mathcal{S}$ represents a problem instance, comprising the source program $C$, its baseline assembly $P$, and the associated test cases $\mathcal{T}$. An action $a\in\mathcal{A}$ corresponds to generating a candidate assembly program $\tilde{P}$. The reward function $r(s,a)$ evaluates the quality of the generated program based on correctness and performance. We will describe different designs of the reward function later. A policy $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ maps a context $s$ to a probability distribution over actions and samples an action $a\in\mathcal{A}$ stochastically. Given a distribution $\mu$ over problem instances, the expected performance of a policy $\pi$ under reward function $r$ is expressed as $\mathbb{E}_{s\sim\mu,a\sim\pi(\cdot\mid s)}\left[r(s,a)\right]$. The objective is to find a policy that maximizes this expected reward.
Optimization with PPO.
We train the policy using Proximal Policy Optimization (PPO) [58], a first-order policy-gradient algorithm that stabilizes training by constraining each policy update to remain close to the previous one. Specifically, PPO maximizes a clipped surrogate objective of the form $\mathbb{E}_{s,a}\left[\min\left(\rho(\theta)\hat{A},\;\operatorname{clip}(\rho(\theta),1-\epsilon,1+\epsilon)\,\hat{A}\right)\right],$ where $\rho(\theta)=\pi_{\theta}(a\mid s)/\pi_{\theta_{\text{old}}}(a\mid s)$ is the probability ratio between the current and previous policy, $\hat{A}$ is the estimated advantage of action $a$ in state $s$, and $\epsilon$ is a clipping coefficient that limits the policy update to a small trust region. We use a critic model to estimate $\hat{A}$, and compute rewards based on the correctness and execution time of the generated program, eliminating the need for a separate reward model.
Reward Function Design.
As defined in our contextual bandit setup, the reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ assigns a scalar score to each (context, action) pair. Each context $s\in\mathcal{S}$ consists of the source program $C$, the baseline assembly $P$, and a test set $\mathcal{T}=\{(x_{i},y_{i})\}_{i=1}^{n}$. An action $a\in\mathcal{A}$ corresponds to a generation procedure that produces a candidate assembly program $\tilde{P}=\texttt{gen}(a)$.
We define two auxiliary metrics for computing reward:
$\operatorname{pass}(s,a)=\tfrac{1}{|\mathcal{T}|}\sum_{(x,y)\in\mathcal{T}}\mathbf{1}[\tilde{P}(x)=y],\quad\operatorname{speedup}(s,a)=t(P)/t(\tilde{P}),$
which respectively denote the fraction of test cases passed and the speedup of the generated program $\tilde{P}$ relative to the baseline $P$. We evaluate two reward function variants:
  1. 1.
    Correctness-Guided Speedup (CGS):
    $r(s,a)=\begin{cases}-1,&\text{if }\tilde{P}\text{ fails to compile},\\[4.0pt] \operatorname{pass}(s,a),&\text{if some tests fail},\\[4.0pt] 1+\alpha\cdot\operatorname{speedup}(s,a),&\text{if all tests pass}.\end{cases}$
  2. 2.
    Speedup-Only (SO):
    $r(s,a)=\begin{cases}0,&\text{if }\tilde{P}\text{ fails to compile or any test fails},\\[4.0pt] \operatorname{speedup}(s,a),&\text{otherwise}.\end{cases}$
In CGS, the constant $\alpha$ controls the relative importance of speedup once full correctness is achieved (i.e., all test cases pass). The CGS reward provides a dense signal by assigning intermediate credit for successful compilation and partially correct outputs, guiding the policy even when the final objective is not yet met. In contrast, SO defines a more direct and sparse objective, assigning nonzero reward only to programs that are both correct and performant, thereby rewarding only the terminal goal of achieving speedup.
4 Experimental Setup
Dataset.
Split # Prog. Avg. Tests Avg. LOC
C Assembly
Training 7,872 8.86 22.3 130.3
Evaluation 200 8.92 21.9 133.3
Table 1: Dataset statistics across training and evaluation splits. LOC = lines of code.
We describe our dataset construction approach in Section 3.2. Each instance consists of a C source program $C$, the corresponding gcc -O3 compiled assembly $P$, and a set of test cases $\mathcal{T}$ for correctness evaluation. The final dataset contains 7,872 training programs and 200 evaluation programs, with average program lengths and test case counts summarized in Table 1.
Prompts.
For each instance, we construct a prompt that includes the original C program along with the generated assembly using gcc -O3. All test cases are withheld from the model. The model is instructed to generate only the optimized x86-64 assembly code. We show the prompt template in Section A.2.
Metrics.
We evaluate each model using both correctness and performance metrics. Compile pass is the percentage of problems for which the generated assembly compiles to binary executable successfully, while test pass is the percentage of problems where the compiled code passes all test cases. For a given problem, any single failed test case is considered a failure for the test pass metric. Both metrics are computed across the entire validation set. For performance, we measure the relative speedup over the gcc -O3 baseline. As defined in Section 3.1, we assign a default speedup of 1× to any candidate that fails to compile, fails any test case, or is slower than the baseline. This reflects the practical setting where a system can fall back to the gcc -O3 output, resulting in no performance gain. We report the 25th, 50th (median), and 75th percentiles of speedup to capture distributional behavior, along with the average speedup over the entire evaluation set.
Models.
We evaluate 21 state-of-the-art language models spanning a diverse range of architectures. Our benchmark includes frontier proprietary models such as gpt-4o [64], o4-mini, gemini-2.0-flash-001 [65], and claude-3.7-sonnet, as well as open-source families such as Llama [66], DeepSeek [67], and Qwen [25]. In addition, we include models distilled from DeepSeek-R1 [68] based on Qwen and Llama. Finally, we evaluate recent compiler foundation models [39, 40] that are pre-trained on assembly code, building upon Code Llama and designed specifically for compiler-related tasks. All open-source models are instruction-tuned.
Performance Measurement.
To ensure an accurate performance evaluation, we use hyperfine [69], a benchmarking tool that reduces measurement noise by performing warmup runs followed by repeated timed executions. For each program’s execution, we discard the first three runs and report the average runtime over the next ten runs.
Implementation.
We implement our customized reinforcement learning reward functions within the verl framework [70], which enables fine-tuning of LLMs using Proximal Policy Optimization (PPO). As part of this setup, we build a task-specific environment that handles program compilation, test execution, and runtime measurement, as detailed in Section 3.3. This environment provides the model with direct scalar feedback based on both functional correctness and execution performance.
Training Configurations.
Among all evaluated models (see Table 2), we select Qwen2.5-Coder-7B-Instruct for training due to its strongest correctness results and substantial room for performance improvement, while intentionally avoiding compiler-specific foundation models to preserve generality. Training is performed on a single node with four A100 GPUs. Full hyperparameter settings are provided in Table A1.
5 Results
5.1 Evaluation of Different Models
Model Compile Pass Test Pass Speedup Percentiles Average Speedup
25th 50th 75th
DS-R1-Distill-Qwen-1.5B 0.0% 0.0% 1.00× 1.00× 1.00× 1.00×
DeepSeek-R1 0.0% 0.0% 1.00× 1.00× 1.00× 1.00×
DS-R1-Distill-Llama-70B 5.5% 0.0% 1.00× 1.00× 1.00× 1.00×
DS-R1-Distill-Qwen-14B 11.5% 0.5% 1.00× 1.00× 1.00× 1.00×
gpt-4o-mini 44.5% 1.0% 1.00× 1.00× 1.00× 1.00×
Llama-4-Maverick-17B 77.5% 7.0% 1.00× 1.00× 1.00× 1.02×
Llama-3.2-11B 84.0% 21.0% 1.00× 1.00× 1.00× 1.02×
gpt-4o 81.0% 5.0% 1.00× 1.00× 1.00× 1.02×
Llama-4-Scout-17B 68.5% 5.5% 1.00× 1.00× 1.00× 1.02×
o4-mini 25.0% 4.5% 1.00× 1.00× 1.00× 1.02×
gemini-2.0-flash-001 57.5% 4.0% 1.00× 1.00× 1.00× 1.03×
Qwen2.5-72B 59.5% 7.5% 1.00× 1.00× 1.00× 1.03×
Llama-3.2-90B 82.5% 15.0% 1.00× 1.00× 1.00× 1.05×
Qwen2.5-Coder-7B 79.0% 61.0% 1.00× 1.00× 1.00× 1.10×
DeepSeek-V3 94.0% 43.0% 1.00× 1.00× 1.40× 1.21×
claude-3.7-sonnet 94.5% 58.5% 1.00× 1.10× 1.45× 1.22×
llm-compiler-7b-ftd 2.0% 2.0% 1.00× 1.00× 1.00× 1.00×
llm-compiler-13b-ftd 2.5% 2.0% 1.00× 1.00× 1.00× 1.01×
llm-compiler-7b 55.0% 54.0% 1.00× 1.00× 1.00× 1.09×
llm-compiler-13b 60.5% 59.5% 1.00× 1.27× 1.63× 1.34×
Qwen2.5-Coder-7B-PPO (Ours) 96.0% 96.0% 1.21× 1.42× 1.66× 1.47×
Table 2: Comparison of LLMs on our assembly optimization benchmark. We report compilation success rate, test pass rate, and average speedup over the gcc -O3 baseline. All open-source models are instruction-tuned. We evaluate general-purpose foundational models, compiler-specific foundation models, and our PPO-trained model, which improves average speedup from 1.10× to 1.47×.
Table 2 presents results across evaluated models. Most models struggle to generate performant assembly: the majority yield only 1.00× speedup, with low compile and test pass rates. Among baseline models, claude-3.7-sonnet and DeepSeek-V3 perform best, achieving test pass rates above 40% and average speedups of 1.22× and 1.21×, respectively. Notably, some models such as DeepSeek-R1 fail to generate any valid assembly, and o4-mini achieves only 4.5% test pass. These results underscore the difficulty of the task and motivate the need for a task-specific approach.
Compiler foundation models (prefixed with llm-compiler-) are pretrained on assembly code and compiler intermediate representations. Among them, llm-compiler-13b demonstrates strong performance in both correctness and speedup. In contrast, the fine-tuned variants (-ftd) perform poorly, likely because they are adapted for tasks such as disassembling x86-64 and ARM assembly into LLVM-IR, rather than optimizing assembly code for execution performance.
We select Qwen2.5-Coder-7B-Instruct for RL training due to its strong compile pass rate (79.0%) and highest test pass rate (61.0%) among models. After PPO fine-tuning, it achieves 96.0% on both metrics and increases average speedup from 1.10× to 1.47×. Notably, it is the only model to exhibit meaningful speedup even at the 25th percentile, and it outperforms all other models across all evaluation metrics, including correctness, average speedup, and speedup percentiles.
Method Compile Pass Test Pass Avg. Speedup
Base Model 79.0% 61.0% 1.10x
RL w/ CGS 95.5% 94.5% 1.38×
RL w/ SO 96.0% 96.0% 1.47×
Table 3: Ablation study comparing reward function variants. CGS provides intermediate reward shaping, while SO uses a sparse and terminal signal.
5.2 Ablation Study of Reward Function Design
We evaluate two reward designs for RL: Correctness-Guided Speedup (CGS) and Speedup-Only (SO). CGS penalizes compilation failures, rewards partial correctness, and scales final reward by speedup once all tests pass. SO uses speedup as the sole reward, but only when all tests pass.
As shown in Table 3, both variants achieve high compile pass rates and test pass rates, but SO yields better performance. Removing intermediate shaping appears to help the model focus on terminal objectives. We also tried varying the CGS scaling factor $\alpha$ (5 or 10) and found that it has a negligible effect.
These results indicate that sparse, terminal rewards (SO) are more effective in this setting. Since the base model already reaches 61.0% test pass, correctness is not the bottleneck; optimizing directly for speedup offers a stronger training signal.
Method Compile Pass Test Pass Avg. Speedup
w/ O3 w/o O3 w/ O3 w/o O3 w/ O3 w/o O3
DeepSeek-V3 94.0% 25.5% 43.0% 4.5% 1.21× 1.02×
claude-3.7-sonnet 94.5% 53.0% 58.5% 16.0% 1.22× 1.07×
llm-compiler-7b 55.0% 12.0% 54.0% 0.0% 1.09× 1.00×
llm-comiler-13b 60.5% 2.0% 59.5% 0.5% 1.34× 1.00×
Qwen2.5-Coder-7B 79.0% 0.0% 61.0% 0.0% 1.10× 1.00×
Qwen2.5-Coder-7B-PPO 96.0% 0.0% 96.0% 0.0% 1.47× 1.00×
Table 4: Ablation study on the impact of including gcc -O3 baseline assembly in the prompt. Each metric is reported with and without access to the baseline assembly generated by the compiler.
5.3 Can LLMs Directly Compile Programs without Baseline Assembly?
In our main evaluation, we always include the baseline assembly generated by gcc -O3 in the prompt. While this baseline offers a strong starting point for further optimization, it may also bias the model toward replicating patterns from the compiler’s output. In this subsection, we investigate a more challenging setup: Can large language models directly compile C code into performant assembly without relying on the compiler-generated baseline?
We compare two settings in evaluation: (1) providing both the C source and the gcc -O3 assembly (default, “w/ O3” in Table 4), and (2) providing only the C source (“w/o O3”). For each model, we report compile success rate, test pass rate, and average speedup.
Table 4 shows that removing the baseline assembly leads to severe degradation. For instance, Qwen2.5-Coder-7B-PPO drops from 96.0% correctness and 1.47× speedup to 0.0% and 1.00×, respectively. Even strong models like Claude-3.7-sonnet suffer substantial declines.
These results suggest that direct compilation from C to optimized assembly remains challenging for current LLMs. The compiler output provides a reliable reference for LLMs. This supports our framework design: using gcc -O3 as an effective starting point for reinforcement learning. While future work may explore direct generation from C, we expect it to be substantially more challenging due to the absence of compiler guidance.
5.4 Case Study
Figure 2 presents a representative example where large language models (LLMs), including gpt-4o and claude-3.7-sonnet, discover an optimization that outperforms a state-of-the-art compiler. The original C function computes the population count (i.e., the number of set bits) by repeatedly shifting the input and accumulating its least significant bit. The assembly code produced by gcc -O3 preserves this loop structure, relying on explicit bitwise operations and conditional branches to compute the result.
In contrast, the LLM generates a significantly more concise and efficient implementation that replaces the entire loop with a single popcnt instruction. This instruction, supported by modern x86-64 architectures, performs the same computation in one operation, thereby reducing both instruction count and runtime overhead.
Such a transformation is beyond the reach of gcc -O3, which applies a predetermined sequence of rule-based optimization passes and does not conduct semantic-level rewrites of this kind. In this case, the language model is able to synthesize functionally equivalent code that exploits hardware-level instructions not utilized by the compiler. This demonstrates the potential of language models to optimize assembly by exploring a broader space of semantics-preserving program transformations.
C Code
int f(unsigned long x)
int res = 0;
while (x > 0)
res += x & 1;
x >>= 1;
return res;
GCC -O3 Output
.L0:
xorl %eax, %eax
testq %rdi, %rdi
je .L2
.L1:
movq %rdi, %rdx
andl \$0x1, %edx
addq %rdx, %rax
shrq \$0x1, %rdi
jne .L1
retq
.L2:
retq
LLM Generated
.L0:
popcnt %rdi, %rax
retq
Figure 2: Case study comparing the C code, baseline assembly produced by gcc -O3, and optimized assembly generated by the LLM. The model successfully replaces the loop with the specialized hardware instruction popcnt, resulting in a significantly more concise implementation.
6 Discussion
Limitations.
A key limitation of our approach is the absence of formal correctness guarantees. Although we validate generated programs using input-output test cases, such testing is inherently incomplete and may overlook edge cases. Consequently, unlike fully verified systems such as Stoke [11], LLM-generated assembly may produce wrong results or exhibit undefined behavior. This limitation reflects a broader challenge in programming languages: verifying the semantic equivalence between two arbitrary programs is undecidable in the general case.
Another limitation lies in the inherent randomness of performance measurement on real hardware. Although we mitigate noise through repeated measurements, low-level hardware fluctuations can still introduce variability into speedup estimates. Such nondeterminism is difficult to eliminate entirely in real-world settings. While prior work has adopted simulator-based evaluation [6], simulators may fail to faithfully capture the actual hardware performance.
Finally, the observed performance gains may not generalize across machines. The model may implicitly learn and exploit hardware-specific characteristics like cache size. As a result, a model trained on one machine may not retain its effectiveness when deployed on a different machine.
Future Work.
While we use Proximal Policy Optimization (PPO), future work may explore alternative reinforcement learning algorithms such as GRPO [71]. Expanding to larger and more diverse datasets, particularly those involving performance-critical code beyond competitive programming, would make the setting more realistic and applicable. Combining reinforcement learning with supervised fine-tuning may also be beneficial, although it remains unclear whether training on gcc -O3 outputs would provide additional gains. Another direction is to incorporate an interactive refinement loop, where the model iteratively updates its output using feedback from errors or performance measurements. Finally, extending our approach from x86-64 to other architectures such as MIPS, ARM, or GPU programming could broaden its applicability and impact.
7 Conclusion
We explore the use of large language models (LLMs) for optimizing assembly code, a setting where fine-grained control over execution enables performance improvements that are difficult to express in high-level languages. While traditional compilers rely on fixed rule-based transformations, they face fundamental limitations due to the complexity of the optimization space. To address this, we apply reinforcement learning to fine-tune LLMs with Proximal Policy Optimization (PPO), using a reward function based on correctness and speedup over the gcc -O3 baseline. To support this effort, we introduce a benchmark of 8,072 real-world C programs with compiler-generated baseline assembly and test cases. Our resulting model, Qwen2.5-Coder-7B-PPO, achieves the highest compile and test pass rates (96.0%) and the best average speedup (1.47×), outperforming all 20 other models evaluated across all metrics. These results indicate that reinforcement learning can unlock the potential of LLMs to serve as effective optimizers for assembly code performance.
Appendix A Appendix
A.1 Training Configurations
Component Setting
Base model Qwen2.5-Coder-7B-Instruct
Actor’s learning rate 1e-6
Critic’s learning rate 1e-5
Batch size 16
Epoch 1
Max prompt length 2000 tokens
Max response length 2000 tokens
Gradient checkpointing Enabled (both actor and critic)
Rollout temperature 0.5
Hardware 4× A100 GPUs
Table A1: Key training configurations for VERL PPO fine-tuning.
A.2 Prompt Template
Prompt Template Given the following C code and assembly code, your task is to generate highly optimized x86-64 assembly code.
C Code:
<C code here>
Assembly Code:
<baseline assembly code here produced by gcc -O3>
Only output the optimized assembly code. Do not include any other text. Do not write any comments in the assembly code. Wrap the assembly code in assembly tags. Optimized Assembly Code:

### v2

@@
-Improving Assembly Code Performance with Large Language Models via Reinforcement Learning
+SuperCoder: Assembly Program Superoptimization with Large Language Models
 Abstract
-Large language models (LLMs) have demonstrated strong performance across a wide range of programming tasks, yet their potential for code optimization remains underexplored.
-This work investigates whether LLMs can optimize the performance of assembly code, where fine-grained control over execution enables improvements that are difficult to express in high-level languages.
-We present a reinforcement learning framework that trains LLMs using Proximal Policy Optimization (PPO), guided by a reward function that considers both functional correctness, validated through test cases, and execution performance relative to the industry-standard compiler gcc -O3.
-To support this study, we introduce a benchmark of 8,072 real-world programs.
-Our model, Qwen2.5-Coder-7B-PPO, achieves 96.0% test pass rates and an average speedup of 1.47× over the gcc -O3 baseline, outperforming all 20 other models evaluated, including Claude-3.7-sonnet.
-These results indicate that reinforcement learning can unlock the potential of LLMs to serve as effective optimizers for assembly code performance.
+Superoptimization is the task of transforming a program into a faster one while preserving its input–output behavior.
+In this work, we investigate whether large language models (LLMs) can serve as superoptimizers, generating assembly programs that outperform code already optimized by industry-standard compilers.
+We construct the first large-scale benchmark for this problem, consisting of 8,072 assembly programs averaging 130 lines, in contrast to prior datasets restricted to 2–15 straight-line, loop-free programs.
+We evaluate 23 LLMs on this benchmark and find that the strongest baseline, Claude-opus-4, achieves a 51.5% test-passing rate and a 1.43× average speedup over gcc -O3.
+To further enhance performance, we fine-tune models with reinforcement learning, optimizing a reward function that integrates correctness and performance speedup.
+Starting from Qwen2.5-Coder-7B-Instruct (61.4% correctness, 1.10× speedup), the fine-tuned model SuperCoder attains 95.0% correctness and 1.46× average speedup.
+Our results demonstrate, for the first time, that LLMs can be applied as superoptimizers for assembly programs, establishing a foundation for future research in program performance optimization beyond compiler heuristics.
 1 Introduction
+Superoptimization is the task of transforming a program into a faster one while preserving its input-output behavior.
+In this work, we investigate whether large language models (LLMs) can perform superoptimization by generating assembly code that surpasses the performance of compiler outputs.
+Decades of research have tackled the problem of code optimization, giving rise to two main approaches.
+The first develops better algorithms for rule-based transformations in compilers (Wolf & Lam, 1991).
+However, given the vast space of possible transformations, compiler-optimized code is not guaranteed to be optimal and often leaves performance untapped (Center et al., 1971; Theodoridis et al., 2022).
+The second, superoptimization, develops search algorithms that directly explore the space of all possible programs, aiming to discover the correct variant with the best performance rather than relying on a fixed set of transformation rules (Schkufza et al., 2013).
+Superoptimization is an aggressive form of program optimization that can outperform compiler-optimized code, yet existing literature has focused on very short, straight-line assembly programs without loops.
+Prior work has primarily relied on CPU-based search heuristics, which fail to scale to larger programs (Schkufza et al., 2013; Phothilimthana et al., 2016; Koenig et al., 2021b); available datasets include at most 15 lines of straight-line assembly (Koenig et al., 2021a).
 Figure 1: Overview of the assembly code optimization task.
-Given a C program and its baseline assembly from gcc -O3, an LLM is fine-tuned with Proximal Policy Optimization (PPO) to generate improved assembly.
+Given a C program and its baseline assembly from gcc -O3, an LLM is fine-tuned with PPO or GRPO to generate improved assembly.
 The reward function reflects correctness and performance based on test execution.
-Recent advances in large language models (LLMs) have achieved state-of-the-art solutions across a wide range of programming tasks [1, 2, 3, 4, 5].
-However, their potential for program optimization remains underexplored.
-Generating highly optimized code is critical in performance-sensitive domains, and prior work has investigated the use of LLMs to optimize C++ and Python programs [6, 7, 8].
-In this work, we aim to utilize LLMs to improve the performance of assembly code, extending their capabilities beyond optimization for high-level languages.
-Assembly code optimization is traditionally the responsibility of compilers.
-While modern compilers apply a series of rule-based transformations to improve performance, such a design introduces the classic phase ordering problem [9], where the order of optimizations can substantially affect the performance of the generated code.
-Due to the inherent complexity of the optimization task, especially the vast space of possible transformation sequences, compilers face fundamental challenges in converging to optimal code, often leaving significant performance on the table [10].
-An alternative approach is superoptimization, which searches the space of all programs that are functionally equivalent to the compiler’s output, aiming to identify the most performant variant.
-In principle, this strategy may yield optimal code.
-However, the search space grows exponentially with program size, making exhaustive exploration computationally infeasible in practice.
-Furthermore, prior work on superoptimization [11, 12] has primarily targeted loop-free, straight-line code, where it is more tractable to formally verify that the optimized code is semantically equivalent to the original.
-As a result, these approaches are not directly applicable to most real-world programs with loops.
-In this work, we explore using large language models to optimize the performance of assembly code.
-Compared with high-level languages such as Python or C++, assembly code operates closer to the hardware, offering fine-grained control over execution and enabling optimizations that are difficult to express or realize in higher-level code.
-However, this setting poses several challenges.
-Assembly code is relatively rare and may be underrepresented in pretraining corpora [13], making it harder for LLMs to reason effectively about their behavior.
-Furthermore, industry compilers such as GCC have been extensively tuned by performance engineers over decades.
-Achieving additional speedups beyond gcc -O3 (the compiler’s highest optimization level) is a technically challenging task.
-To address the challenges of low-level code optimization, we apply reinforcement learning to enhance the ability of LLMs to optimize assembly code.
-As shown in Figure 1, we use Proximal Policy Optimization (PPO) to train an LLM using a reward function that considers both correctness and performance.
-Correctness is evaluated based on whether the generated code passes program-specific test cases, and performance is measured by its speedup relative to the baseline produced by gcc -O3.
-To support this setting, we construct a new dataset of 8,072 assembly programs derived from real-world competitive programming submissions.
-Each instance includes input-output test cases and baseline assembly code generated by the compiler at the highest optimization level, which serves as the starting point for further optimization.
-We evaluate our approach on the proposed benchmark and find that reinforcement learning substantially improves the ability of LLMs to optimize assembly code.
-Starting from the base model Qwen2.5-Coder-7B-Instruct, which achieves a modest 1.10x speedup over the gcc -O3 baseline, our PPO-trained model reaches 1.47× average speedup and improves both compile and test pass rates to 96.0%.
-It achieves the strongest performance across all evaluation metrics, outperforming all 20 other models evaluated, including Claude-3.7-sonnet.
-Ablation studies show that reward functions emphasizing final speedup, rather than intermediate correctness signals, lead to more effective training.
+In this work, we explore the use of LLMs as a superoptimizer to improve the performance of assembly code.
+In contrast to most prior work on code generation from natural language (Chen et al., 2021; Austin et al., 2021; Hendrycks et al., 2021; Zhuo et al., 2024), we tackle a fundamentally different and more technically demanding task: improving assembly code that has already been optimized by the industry-standard compiler at its highest optimization level (gcc -O3).
+Compilers have been refined over decades of expert-driven development, and surpassing them remains a central challenge in programming languages, as compilers form the foundation of all software.
+Unlike high-level programming languages (e.g., Python or C), large-scale, high-quality assembly datasets are scarce.
+As the first study in this direction, we construct a dataset of 8,072 assembly programs.
+Each instance includes input–output test cases and baseline assembly generated by the compiler at its highest optimization level (gcc -O3), which serves as the starting point for further optimization.
+In contrast, the datasets commonly used in the superoptimization community (Warren, 2013; Schkufza et al., 2013; Koenig et al., 2021a) are either extremely limited in size, containing only 25 programs, or consist of toy examples with 2 to 15 instructions, 74% of which have between 3 and 6 instructions and contain no loops.
+Our dataset is substantially larger, with assembly programs averaging 130 lines and including loops.
+Our test suites achieve 96.2% line and 87.3% branch coverage, demonstrating strong test quality.
+Our dataset represents a substantial step forward in scale for evaluating superoptimization techniques.
+Beyond evaluating existing models, we also apply reinforcement learning (RL) for fine-tuning to further enhance their capabilities.
+We use widely adopted algorithms, including Proximal Policy Optimization (PPO) and Group Relative Policy Optimization (GRPO), to train an LLM with a reward function that integrates both correctness and performance speedup.
+Prior work on LLM-based performance optimization has explored alternative methodologies such as supervised fine-tuning (Shypula et al., 2023), chain-of-thought prompting (Liu et al., 2024c), agent-based frameworks (Wei et al., 2025b; c), and preference learning (Du et al., 2024).
+Our approach optimizes speedup explicitly in the reward function, making reinforcement learning well suited to superoptimization.
+To our knowledge, this is the first application of reward-based RL to LLMs for code performance optimization, with correctness and speedup jointly encoded in the objective.
+We evaluate 23 LLMs on this task and find that the best-performing model, Claude-opus-4, achieves a 51.5% test-passing rate and an average speedup of 1.43× over the compiler-optimized baseline (gcc -O3).
+Our reinforcement learning approach is highly effective: starting from the base model Qwen2.5-Coder-7B-Instruct, which achieves 61.4% correctness and a modest 1.10× speedup, the fine-tuned model SuperCoder attains 95.0% correctness and 1.46× average speedup.
+We further analyze the code transformation patterns learned by the model.
 In summary, our contributions are as follows:
-  • We introduce the task of optimizing assembly code performance using large language models, aiming for fine-grained performance improvements beyond what traditional compiler optimizations can achieve.
-  • We construct a dataset of real-world C programs paired with the corresponding assembly code generated by the gcc -O3 baseline.
-Using this dataset, we explore improving LLMs on this task through reinforcement learning, applying Proximal Policy Optimization (PPO).
-  • We evaluate 21 LLMs on the proposed benchmark and show that our training substantially improves performance: Qwen2.5-Coder-7B-PPO achieves the highest compile and test pass rates, as well as the best average speedup (1.47×) over the gcc -O3 baseline, outperforming all other models (including Claude-3.7-sonnet) across all evaluation metrics.
+  • We are the first to introduce superoptimization as a task for LLMs, a technically demanding challenge that aims to improve assembly code already optimized by industry-standard compilers.
+  • We construct the first large-scale dataset of 8,072 assembly programs, averaging 130 lines.
+This far surpasses prior loop-free datasets under 15 lines and marks a substantial step forward in scale and realism for evaluating superoptimization techniques.
+  • We evaluate 23 LLMs on the benchmark and show that RL-based training substantially improves performance: fine-tuning Qwen2.5-Coder-7B-Instruct (61.4% correctness, 1.10× speedup) results in SuperCoder with 95.0% correctness and 1.46× speedup.
 2 Related Work
@@
 Benchmarks for evaluating large language models (LLMs) on code generation from natural language specifications have received increasing attention.
-Notable examples include HumanEval [1], MBPP [2], APPS [3], and more recent efforts [14, 15, 16, 17].
-In parallel, many models have been developed to enhance code generation capabilities, such as Codex [1], AlphaCode [18], CodeGen [19], InCoder [20], StarCoder [21], DeepSeek-Coder [22], Code Llama [23], and others [24, 25].
-Beyond code generation, LLMs have been applied to real-world software engineering tasks including automated program repair [26, 27], software testing [28, 29], bug localization [30], and transpilation [31, 32].
-SWE-bench [4] integrates these tasks into a benchmark for resolving real GitHub issues.
-Building on this, SWE-agent [5] and subsequent works [33, 34] employ an agent-based framework that leverages LLMs to improve the issue resolution process.
+Notable examples include HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), APPS (Hendrycks et al., 2021), and more recent efforts (Liu et al., 2023b; Li et al., 2024; Xia et al., 2024b; Zhuo et al., 2024).
+In parallel, many models have been developed to enhance code generation capabilities, such as Codex (Chen et al., 2021), AlphaCode (Li et al., 2022), CodeGen (Nijkamp et al., 2022), InCoder (Fried et al., 2022), StarCoder (Li et al., 2023), DeepSeek-Coder (Guo et al., 2024), Code Llama (Roziere et al., 2023), and others (Hui et al., 2024; Wei et al., 2025e).
+Beyond code generation, LLMs have been applied to real-world software engineering tasks including automated program repair (Xia & Zhang, 2022; Xia et al., 2023b), software testing (Xia et al., 2023a; Deng et al., 2024), bug localization (Yang et al., 2024a), transpilation (Yang et al., 2024c; Bhatia et al., 2024), equivalence checking (Wei et al., 2025a), and synthesis (Wei et al., 2025d).
+SWE-bench (Jimenez et al., 2023) integrates these tasks into a benchmark for resolving real GitHub issues.
+Building on this, SWE-agent (Yang et al., 2024b) and subsequent works (Xia et al., 2024a; Wei et al., 2025f) employ an agent-based framework that leverages LLMs to improve the issue resolution process.
 Recent work has also explored LLMs for improving program performance.
-CodeRosetta [35] targets automatic parallelization, such as translating C++ to CUDA.
-Other efforts focus on optimizing Python code for efficiency [7, 8] or enabling self-adaptation [36], and improving C++ performance [6].
-Of particular relevance are approaches to low-level code optimization [37, 38].
-The LLM Compiler foundation models [39, 40] are primarily designed for code size reduction and binary disassembly, whereas our work focuses on optimizing assembly code for performance.
-LLM-Vectorizer [41] offers a formally verified solution for auto-vectorization, a specific compiler pass.
+CodeRosetta (TehraniJamsaz et al., 2024) targets automatic parallelization, such as translating C++ to CUDA.
+Other efforts focus on optimizing Python code for efficiency (Du et al., 2024; Liu et al., 2024c) or enabling self-adaptation (Huang et al., 2024), and improving C++ performance (Shypula et al., 2023).
+Of particular relevance are approaches to low-level code optimization (Wei et al., 2024; Ouyang et al., 2025).
+The LLM Compiler foundation models (Cummins et al., 2024; 2025) are primarily designed for code size reduction and binary disassembly, whereas our work focuses on optimizing assembly code for performance.
+LLM-Vectorizer (Taneja et al., 2025) offers a formally verified solution for auto-vectorization, a specific compiler pass.
 In contrast, our work does not restrict the optimization type and uses test-case validation.
@@
 A classic challenge in compilers is the phase-ordering problem, where performance depends heavily on the sequence of optimization passes.
-AutoPhase [42] uses deep reinforcement learning to tackle this, while Coreset [43] employs graph neural networks (GNNs) to guide optimization decisions.
+AutoPhase (Haj-Ali et al., 2020) uses deep reinforcement learning to tackle this, while Coreset (Liang et al., 2023) employs graph neural networks (GNNs) to guide optimization decisions.
 Modern compilers apply extensive rewrite rules but offer no guarantee of optimality.
 Superoptimization seeks the most efficient program among all semantically equivalent variants of the compiler output.
-Traditional methods use stochastic search, such as Markov Chain Monte Carlo [11], with follow-up work improving scalability [44, 12] and extending to broader domains [45, 46].
+Traditional methods use stochastic search, such as MCMC sampling (Schkufza et al., 2013), with follow-up work improving scalability (Phothilimthana et al., 2016; Bunel et al., 2016) and extending to broader domains (Sharma et al., 2015; Churchill et al., 2017).
 These rely on formal verification for correctness, restricting them to small, loop-free programs.
@@
 With the rise of deep learning, substantial attention has turned to optimizing GPU kernel code.
-AutoTVM [47] pioneered statistical cost model-based search for CUDA code optimization, followed by methods such as Ansor [48], AMOS [49], and other recent systems [50, 51, 52].
-More recently, using LLMs as code optimizers has gained popularity [6, 53, 37], with growing attention to reinforcement learning approaches that guide LLMs through reward-based feedback [54, 34].
-CodeRL [55] incorporates unit test-based rewards within an actor-critic framework [56], while PPOCoder [57] extends this with Proximal Policy Optimization (PPO) [58], along with other variants [59].
-Subsequent efforts have adapted RL-based techniques [60] to additional low-resource programming languages, including Verilog [61].
-To the best of our knowledge, our work is the first to apply reinforcement learning to optimize assembly code using LLMs.
-Assembly code offers fine-grained control and potential for significant performance gains, but it remains underexplored due to limited training data and the complexity of low-level semantics.
+AutoTVM (Chen et al., 2018) pioneered statistical cost model-based search for CUDA code optimization, followed by methods such as Ansor (Zheng et al., 2020), AMOS (Zheng et al., 2022), and others (Shao et al., 2022; Zhao et al., 2024; Wu et al., 2024).
+More recently, LLMs have been explored as code optimizers (Shypula et al., 2023; Grubisic et al., 2024; Wei et al., 2024; 2025b; 2025c), with growing interest in reinforcement learning that guides generation through reward signals (Dou et al., 2024; Wei et al., 2025f).
+Rewards are often derived from unit-test correctness (Le et al., 2022; Shojaee et al., 2023; Liu et al., 2023a) or binary preference signals (Liu et al., 2024b; Du et al., 2024).
+To our knowledge, this is the first work to apply reinforcement learning to optimize code performance with LLMs, with concurrent efforts exploring CUDA kernel optimization (Li et al., 2025; Baronio et al., 2025).
 3 Methodology
@@
 Let $C$ be a program written in a high-level language such as C.
-A modern compiler like gcc can compile $C$ into an x86-64 assembly program $P=\texttt{gcc}(C)$, which can then be further assembled into an executable binary.
+A modern compiler like gcc can compile $C$ into an x86-64 assembly program $P=gcc(C)$, which can then be further assembled into an executable binary.
 The assembly program $P$ serves as an intermediate representation that exposes low-level optimization opportunities, making it suitable for aggressive performance improvement.
@@
 Let $\texttt{valid}(P^{\prime})\in\{\texttt{True},\texttt{False}\}$ denote this property.
-We define the set of correct programs as:
+Based on all that we said above, we define the set of correct programs as:
 $\mathcal{S}(P)=\left\{P^{\prime}\mid\texttt{valid}(P^{\prime})\;\land\;\forall(x_{i},y_{i})\in\mathcal{T},\;P^{\prime}(x_{i})=y_{i}\right\}.$
@@
 Let $t(P)$ denote the execution time of $P$ on the test set $\mathcal{T}$, and let $t(P^{\prime})$ be the corresponding execution time for $P^{\prime}$.
-The speedup of $P^{\prime}$ relative to $P$ is defined as:
+The speedup of $P^{\prime}$ relative to $P$ is defined as follows.
+If the LLM-generated program is invalid or slower, we fall back to the baseline and assign a speedup of 1.
 $\text{Speedup}(P^{\prime})=\begin{cases}\frac{t(P)}{t(P^{\prime})}&\text{if }P^{\prime}\in\mathcal{S}(P)\text{ and }t(P^{\prime})<t(P),\\ 1&\text{otherwise}.\end{cases}$
@@
 Although $\mathcal{S}(P)$ captures the correctness criteria, we do not restrict the LLM to generate only valid programs.
-Instead, the model produces arbitrary assembly code, and correctness is verified post hoc via compilation and test execution.
-We train an LLM using reinforcement learning (see Section 3.3) to generate candidates that both satisfy correctness and achieve performance improvements.
+Instead, the model produces arbitrary assembly code, and correctness is validated post hoc via compilation and test execution.
+We train an LLM using reinforcement learning (see Section˜3.3) to generate candidates that both satisfy correctness and achieve performance improvements.
 3.2 Dataset Construction
-We construct our dataset using C programs from CodeNet [62], a large-scale corpus of competitive programming submissions.
-Each dataset instance is a tuple $(C,P,\mathcal{T})$, where $C$ is the original C source code, $P=\texttt{gcc\_O3}(C)$ is the corresponding x86-64 assembly generated by compiling $C$ with gcc at the -O3 optimization level, and $\mathcal{T}=\{(x_{i},y_{i})\}_{i=1}^{n}$ is the test set.
-Since not all CodeNet problems include test inputs, we adopt those provided by prior work [18] to define $x_{i}$, but discard their output labels.
+We construct our dataset using C programs from CodeNet (Puri et al., 2021), a large-scale corpus of competitive programming submissions.
+Each dataset instance is a tuple $(C,P,\mathcal{T})$, where $C$ is the original C source code, $P=gcc(C)$ is the corresponding x86-64 assembly generated by compiling $C$ with gcc at the -O3 optimization level, and $\mathcal{T}=\{(x_{i},y_{i})\}_{i=1}^{n}$ is the test set.
+Since not all CodeNet problems include test inputs, we adopt those provided by prior work (Li et al., 2022) to define $x_{i}$, but discard their output labels.
 Instead, we regenerate each output $y_{i}$ by executing the original submission on input $x_{i}$, as many CodeNet programs are not accepted solutions, and even accepted ones do not reliably pass all test cases.
@@
 To focus on performance-critical cases, we sample programs that exhibit the highest relative speedup from gcc -O0 (no optimization) to gcc -O3 (maximum optimization).
-Such strategy serves two purposes: (1) it favors programs with complex logic that lead to suboptimal performance under -O0 and can be effectively optimized by -O3, and (2) it creates a more challenging setting by starting from code that has already benefited from aggressive compiler optimizations.
+Such a strategy serves two purposes: (1) it favors programs with complex logic that lead to suboptimal performance under -O0 and can be effectively optimized by -O3, and (2) it creates a more challenging setting by starting from code that has already benefited from aggressive compiler optimizations.
 If an LLM can generate code that further improves upon gcc -O3, it suggests that the model can outperform the compiler’s “expert” solution.
-The final dataset consists of 7,872 training programs and 200 held-out evaluation programs, with additional statistics provided in Section 4.
+The final dataset consists of 7,872 training programs and 200 held-out evaluation programs, with additional statistics provided in Section˜4.
 3.3 Reinforcement Learning
-We conceptualize our task as a standard contextual multi-armed bandit problem [63], defined by a context space $\mathcal{S}$, an action space $\mathcal{A}$, and a reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$.
+We conceptualize our task as a standard contextual multi-armed bandit problem (Lu et al., 2010), defined by a context space $\mathcal{S}$, an action space $\mathcal{A}$, and a reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$.
 Each context $s\in\mathcal{S}$ represents a problem instance, comprising the source program $C$, its baseline assembly $P$, and the associated test cases $\mathcal{T}$.
@@
 The objective is to find a policy that maximizes this expected reward.
-Optimization with PPO.
-We train the policy using Proximal Policy Optimization (PPO) [58], a first-order policy-gradient algorithm that stabilizes training by constraining each policy update to remain close to the previous one.
-Specifically, PPO maximizes a clipped surrogate objective of the form $\mathbb{E}_{s,a}\left[\min\left(\rho(\theta)\hat{A},\;\operatorname{clip}(\rho(\theta),1-\epsilon,1+\epsilon)\,\hat{A}\right)\right],$ where $\rho(\theta)=\pi_{\theta}(a\mid s)/\pi_{\theta_{\text{old}}}(a\mid s)$ is the probability ratio between the current and previous policy, $\hat{A}$ is the estimated advantage of action $a$ in state $s$, and $\epsilon$ is a clipping coefficient that limits the policy update to a small trust region.
-We use a critic model to estimate $\hat{A}$, and compute rewards based on the correctness and execution time of the generated program, eliminating the need for a separate reward model.
+Optimization with PPO and GRPO.
+[…]
 Reward Function Design.
@@
 We define two auxiliary metrics for computing reward:
-$\operatorname{pass}(s,a)=\tfrac{1}{|\mathcal{T}|}\sum_{(x,y)\in\mathcal{T}}\mathbf{1}[\tilde{P}(x)=y],\quad\operatorname{speedup}(s,a)=t(P)/t(\tilde{P}),$
-which respectively denote the fraction of test cases passed and the speedup of the generated program $\tilde{P}$ relative to the baseline $P$. We evaluate two reward function variants:
-  1. 1.
-    Correctness-Guided Speedup (CGS):
-    $r(s,a)=\begin{cases}-1,&\text{if }\tilde{P}\text{ fails to compile},\\[4.0pt] \operatorname{pass}(s,a),&\text{if some tests fail},\\[4.0pt] 1+\alpha\cdot\operatorname{speedup}(s,a),&\text{if all tests pass}.\end{cases}$
-  2. 2.
-    Speedup-Only (SO):
-    $r(s,a)=\begin{cases}0,&\text{if }\tilde{P}\text{ fails to compile or any test fails},\\[4.0pt] \operatorname{speedup}(s,a),&\text{otherwise}.\end{cases}$
-In CGS, the constant $\alpha$ controls the relative importance of speedup once full correctness is achieved (i.e., all test cases pass).
-The CGS reward provides a dense signal by assigning intermediate credit for successful compilation and partially correct outputs, guiding the policy even when the final objective is not yet met.
-In contrast, SO defines a more direct and sparse objective, assigning nonzero reward only to programs that are both correct and performant, thereby rewarding only the terminal goal of achieving speedup.
+$t(P~),\operatorname{pass}(s,a)=\tfrac{1}{|\mathcal{T}|}\sum_{(x,y)\in\mathcal{T}}\mathbf{1}[\tilde{P}(x)=y],\quad\operatorname{speedup}(s,a)$$t(\tilde{P})$,
+which denote the fraction of test cases passed and the speedup of the generated program $\tilde{P}$ relative to the baseline $P$. We use the following reward function during training:
+$r(s,a)=\begin{cases}0,&\text{if }\operatorname{pass}(s,a)<1,\\[4.0pt] \operatorname{speedup}(s,a),&\text{if }\operatorname{pass}(s,a)=1.\end{cases}$
+If a generated program fails to compile or does not pass all tests, its reward is set to 0, with no partial credit for partial correctness.
+Only when the code compiles and passes all tests is a positive reward assigned, equal to the achieved speedup.
 4 Experimental Setup
 Dataset.
+We describe our dataset construction approach in Section˜3.2.
+Each instance consists of a C source program $C$, the corresponding gcc -O3 compiled assembly $P$, and a set of test cases $\mathcal{T}$ for correctness evaluation.
+The final dataset contains 7,872 training programs and 200 evaluation programs, with average program lengths and test case counts summarized in Table˜1, and additional analysis below.
+Test Coverage.
+Our dataset includes test cases for every program.
+Rather than relying directly on the original submissions, we re-run each program on its inputs to generate correct outputs, thereby fixing errors in prior datasets.
+The resulting test suites of our evaluation dataset achieve an average of 96.2% line coverage and 87.3% branch coverage, demonstrating high test quality.
 Split # Prog. Avg. Tests Avg. LOC
@@
 Table 1: Dataset statistics across training and evaluation splits. LOC = lines of code.
-We describe our dataset construction approach in Section 3.2.
-Each instance consists of a C source program $C$, the corresponding gcc -O3 compiled assembly $P$, and a set of test cases $\mathcal{T}$ for correctness evaluation.
-The final dataset contains 7,872 training programs and 200 evaluation programs, with average program lengths and test case counts summarized in Table 1.
+Speedup by Compilers.
+We quantify compiler optimizations by comparing gcc -O0 with gcc -O3 on the evaluation dataset and observe a mean speedup of 2.65×.
+This demonstrates the substantial effect of compiler optimizations and confirms that performance improvements in our dataset are measurable.
+Building on this baseline, we investigate whether LLMs can further enhance performance beyond the 2.65× speedup provided by the compiler.
 Prompts.
 For each instance, we construct a prompt that includes the original C program along with the generated assembly using gcc -O3.
-All test cases are withheld from the model.
-The model is instructed to generate only the optimized x86-64 assembly code.
-We show the prompt template in Section A.2.
+Test cases are withheld from the model.
+The model is instructed to generate the optimized assembly code.
+We show the prompt template in Appendix A.3.
 Metrics.
@@
 For performance, we measure the relative speedup over the gcc -O3 baseline.
-As defined in Section 3.1, we assign a default speedup of 1× to any candidate that fails to compile, fails any test case, or is slower than the baseline.
+As defined in Section˜3.1, we assign a default speedup of 1× to any candidate that fails to compile, fails any test case, or is slower than the baseline.
 This reflects the practical setting where a system can fall back to the gcc -O3 output, resulting in no performance gain.
@@
 Models.
-We evaluate 21 state-of-the-art language models spanning a diverse range of architectures.
-Our benchmark includes frontier proprietary models such as gpt-4o [64], o4-mini, gemini-2.0-flash-001 [65], and claude-3.7-sonnet, as well as open-source families such as Llama [66], DeepSeek [67], and Qwen [25].
-In addition, we include models distilled from DeepSeek-R1 [68] based on Qwen and Llama.
-Finally, we evaluate recent compiler foundation models [39, 40] that are pre-trained on assembly code, building upon Code Llama and designed specifically for compiler-related tasks.
+We evaluate 23 state-of-the-art language models spanning a diverse range of architectures.
+Our benchmark includes frontier proprietary models such as gpt-4o (Achiam et al., 2023), o4-mini, gemini-2.0-flash-001 (Team et al., 2023), and claude-3.7-sonnet, as well as open-source families such as Llama (Touvron et al., 2023), DeepSeek (Liu et al., 2024a), and Qwen (Hui et al., 2024).
+In addition, we include models distilled from DeepSeek-R1 (Guo et al., 2025) based on Qwen and Llama.
+Finally, we evaluate recent compiler-oriented foundation models (Cummins et al., 2024; 2025), pretrained on assembly code and derived from Code Llama, with a design focus on compiler tasks (listed as llm-compiler in Table˜2).
 All open-source models are instruction-tuned.
 Performance Measurement.
-To ensure an accurate performance evaluation, we use hyperfine [69], a benchmarking tool that reduces measurement noise by performing warmup runs followed by repeated timed executions.
+To ensure an accurate performance evaluation, we use hyperfine (hyp, 2025), a benchmarking tool that reduces measurement noise by performing warmup runs followed by repeated timed executions.
 For each program’s execution, we discard the first three runs and report the average runtime over the next ten runs.
 Implementation.
-We implement our customized reinforcement learning reward functions within the verl framework [70], which enables fine-tuning of LLMs using Proximal Policy Optimization (PPO).
-As part of this setup, we build a task-specific environment that handles program compilation, test execution, and runtime measurement, as detailed in Section 3.3.
+We implement our customized reinforcement learning reward functions within the verl framework (Sheng et al., 2024), which enables fine-tuning of LLMs using PPO and GRPO.
+As part of this setup, we build a task-specific environment that handles program compilation, test execution, and runtime measurement, as detailed in Section˜3.3.
 This environment provides the model with direct scalar feedback based on both functional correctness and execution performance.
 Training Configurations.
-Among all evaluated models (see Table 2), we select Qwen2.5-Coder-7B-Instruct for training due to its strongest correctness results and substantial room for performance improvement, while intentionally avoiding compiler-specific foundation models to preserve generality.
+Among all evaluated models (see Table˜2), we select Qwen2.5-Coder-7B-Instruct for training due to its strong correctness results and substantial room for performance improvement, while intentionally avoiding compiler-specific foundation models to preserve generality.
 Training is performed on a single node with four A100 GPUs.
-Full hyperparameter settings are provided in Table A1.
+Full hyperparameter settings are provided in Appendix A1.
 5 Results
@@
 Llama-3.2-90B 82.5% 15.0% 1.00× 1.00× 1.00× 1.05×
-Qwen2.5-Coder-7B 79.0% 61.0% 1.00× 1.00× 1.00× 1.10×
+Qwen2.5-Coder-7B 77.9% 61.4% 1.00× 1.00× 1.00× 1.10×
+gpt-5 78.5% 6.0% 1.00× 1.00× 1.00× 1.13×
 DeepSeek-V3 94.0% 43.0% 1.00× 1.00× 1.40× 1.21×
 claude-3.7-sonnet 94.5% 58.5% 1.00× 1.10× 1.45× 1.22×
+claude-sonnet-4 87.0% 37.0% 1.00× 1.00× 1.95× 1.30×
+claude-opus-4 90.0% 51.5% 1.00× 1.58× 2.03× 1.43×
 llm-compiler-7b-ftd 2.0% 2.0% 1.00× 1.00× 1.00× 1.00×
@@
 llm-compiler-13b 60.5% 59.5% 1.00× 1.27× 1.63× 1.34×
-Qwen2.5-Coder-7B-PPO (Ours) 96.0% 96.0% 1.21× 1.42× 1.66× 1.47×
 Table 2: Comparison of LLMs on our assembly optimization benchmark.
@@
 All open-source models are instruction-tuned.
-We evaluate general-purpose foundational models, compiler-specific foundation models, and our PPO-trained model, which improves average speedup from 1.10× to 1.47×.
-Table 2 presents results across evaluated models.
+Main Results.
+Table˜2 reports results across evaluated models.
+Most perform poorly on this task, with only a few demonstrating effectiveness as superoptimizers.
 Most models struggle to generate performant assembly: the majority yield only 1.00× speedup, with low compile and test pass rates.
-Among baseline models, claude-3.7-sonnet and DeepSeek-V3 perform best, achieving test pass rates above 40% and average speedups of 1.22× and 1.21×, respectively.
-Notably, some models such as DeepSeek-R1 fail to generate any valid assembly, and o4-mini achieves only 4.5% test pass.
-These results underscore the difficulty of the task and motivate the need for a task-specific approach.
-Compiler foundation models (prefixed with llm-compiler-) are pretrained on assembly code and compiler intermediate representations.
-Among them, llm-compiler-13b demonstrates strong performance in both correctness and speedup.
-In contrast, the fine-tuned variants (-ftd) perform poorly, likely because they are adapted for tasks such as disassembling x86-64 and ARM assembly into LLVM-IR, rather than optimizing assembly code for execution performance.
-We select Qwen2.5-Coder-7B-Instruct for RL training due to its strong compile pass rate (79.0%) and highest test pass rate (61.0%) among models.
-After PPO fine-tuning, it achieves 96.0% on both metrics and increases average speedup from 1.10× to 1.47×.
-Notably, it is the only model to exhibit meaningful speedup even at the 25th percentile, and it outperforms all other models across all evaluation metrics, including correctness, average speedup, and speedup percentiles.
-Method Compile Pass Test Pass Avg. Speedup
-Base Model 79.0% 61.0% 1.10x
-RL w/ CGS 95.5% 94.5% 1.38×
-RL w/ SO 96.0% 96.0% 1.47×
-Table 3: Ablation study comparing reward function variants. CGS provides intermediate reward shaping, while SO uses a sparse and terminal signal.
-5.2 Ablation Study of Reward Function Design
-We evaluate two reward designs for RL: Correctness-Guided Speedup (CGS) and Speedup-Only (SO).
-CGS penalizes compilation failures, rewards partial correctness, and scales final reward by speedup once all tests pass.
-SO uses speedup as the sole reward, but only when all tests pass.
-As shown in Table 3, both variants achieve high compile pass rates and test pass rates, but SO yields better performance.
-Removing intermediate shaping appears to help the model focus on terminal objectives.
-We also tried varying the CGS scaling factor $\alpha$ (5 or 10) and found that it has a negligible effect.
-These results indicate that sparse, terminal rewards (SO) are more effective in this setting.
-Since the base model already reaches 61.0% test pass, correctness is not the bottleneck; optimizing directly for speedup offers a stronger training signal.
-Method Compile Pass Test Pass Avg. Speedup
-w/ O3 w/o O3 w/ O3 w/o O3 w/ O3 w/o O3
-DeepSeek-V3 94.0% 25.5% 43.0% 4.5% 1.21× 1.02×
-claude-3.7-sonnet 94.5% 53.0% 58.5% 16.0% 1.22× 1.07×
-llm-compiler-7b 55.0% 12.0% 54.0% 0.0% 1.09× 1.00×
-llm-comiler-13b 60.5% 2.0% 59.5% 0.5% 1.34× 1.00×
-Qwen2.5-Coder-7B 79.0% 0.0% 61.0% 0.0% 1.10× 1.00×
-Qwen2.5-Coder-7B-PPO 96.0% 0.0% 96.0% 0.0% 1.47× 1.00×
-Table 4: Ablation study on the impact of including gcc -O3 baseline assembly in the prompt. Each metric is reported with and without access to the baseline assembly generated by the compiler.
-5.3 Can LLMs Directly Compile Programs without Baseline Assembly?
-In our main evaluation, we always include the baseline assembly generated by gcc -O3 in the prompt.
-While this baseline offers a strong starting point for further optimization, it may also bias the model toward replicating patterns from the compiler’s output.
-In this subsection, we investigate a more challenging setup: Can large language models directly compile C code into performant assembly without relying on the compiler-generated baseline?
-We compare two settings in evaluation: (1) providing both the C source and the gcc -O3 assembly (default, “w/ O3” in Table 4), and (2) providing only the C source (“w/o O3”).
-For each model, we report compile success rate, test pass rate, and average speedup.
-Table 4 shows that removing the baseline assembly leads to severe degradation.
-For instance, Qwen2.5-Coder-7B-PPO drops from 96.0% correctness and 1.47× speedup to 0.0% and 1.00×, respectively.
-Even strong models like Claude-3.7-sonnet suffer substantial declines.
-These results suggest that direct compilation from C to optimized assembly remains challenging for current LLMs.
-The compiler output provides a reliable reference for LLMs.
-This supports our framework design: using gcc -O3 as an effective starting point for reinforcement learning.
-While future work may explore direct generation from C, we expect it to be substantially more challenging due to the absence of compiler guidance.
-5.4 Case Study
-Figure 2 presents a representative example where large language models (LLMs), including gpt-4o and claude-3.7-sonnet, discover an optimization that outperforms a state-of-the-art compiler.
+Among all models, claude-opus-4 and claude-sonnet-4 perform best, achieving average speedups of 1.43× and 1.30×, respectively.
+Compiler foundation models (prefixed with llm-compiler-) are pretrained on assembly code and compiler IRs.
+Among them, llm-compiler-13b achieves a notable 1.34× speedup, whereas the fine-tuned variants (-ftd) perform poorly, likely because they were adapted for different tasks such as disassembling x86-64 into LLVM IR.
+These results suggest that while superoptimization is inherently difficult, some LLMs can be effective superoptimizers.
+Failure Modes.
+Interestingly, models that are expected to achieve strong results perform (e.g., DeepSeek-R1, GPT-4o) perform poorly on the task of superoptimization, motivating our analysis of their failure modes.
+We find that DeepSeek-R1 consistently fails to generate valid assembly code, resulting in a 0% compilation rate.
+DeepSeek-R1 often produces verbose analysis instead of executable code, spending its entire output length on reasoning about instruction semantics and potential optimizations without actually implementing them.
+We further analyze the failure modes of GPT-4o, which achieves a high compilation rate (81.0%) but exhibits poor correctness (only 5.0% test pass rate).
+The primary correctness issues are as follows: (1) missing critical directives and safety setup, such as stack canary initialization and .cfi_* metadata, which often lead to runtime crashes; (2) incorrect function call conventions, where repeated system calls like scanf are made without proper argument setup, causing undefined behavior; (3) semantic errors in core computations, including incorrect pointer usage or altered algorithm logic, which produce wrong outputs even when the code runs; and (4) over-simplified stack or register management, resulting in memory errors or invalid control flow.
+In summary, GPT-4o tends to sacrifice correctness in pursuit of optimization: it generates syntactically valid assembly but frequently violates low-level conventions necessary for correct and reliable execution.
+5.2 Effectiveness of RL Training
+Improvement.
+Table˜3 presents the results of RL training, averaged over 5 runs with 95% confidence intervals to provide more statistical confidence in the reported improvements.
+We select Qwen2.5-Coder-7B-Instruct for RL training due to its strong test pass rate (61.4%) among models.
+After RL training with PPO, the fine-tuned model SuperCoder attains 95.0% correctness and improves average speedup from 1.10× to 1.46×.
+Its speedup percentiles are 1.17 ± 0.03× (25th), 1.35 ± 0.04× (50th), and 1.64 ± 0.08× (75th) respectively, outperforming the majority of evaluated models.
+PPO versus GRPO.
+We evaluate both PPO-trained and GRPO-trained models and find their performance to be nearly identical.
+SuperCoder trained with GRPO attains 94.7 ± 0.6% correctness and 1.44 ± 0.07× average speedup, which is comparable to SuperCoder trained with PPO (95.0 ± 0.0% correctness and 1.46 ± 0.12× average speedup).
+Model Compile Pass Test Pass Average Speedup
+[…]
+[…]
+[…]
+Table 3: Performance of the base model and the models fine-tuned with RL. Results include compilation success, test pass rates, and average speedup, reported over 5 runs with 95% confidence intervals.
+5.3 Analysis of Learned Program Transformations
+To better understand why LLMs can further optimize assembly programs already optimized by industry-standard compilers, we analyze all 200 evaluation programs by comparing the gcc -O3 output with the assembly generated by our PPO-trained model SuperCoder.
+We categorize the learned transformations into four types: (1) code layout and instruction scheduling, which account for 98.5% of the changes and include basic block placement, instruction alignment, and reordering to improve cache behavior and hide latencies; (2) register allocation, representing 16% of the changes, where the model selects alternative registers; (3) control flow optimization (1.5%); and (4) instruction selection (1.0%).
+Since a single problem’s transformation may belong to multiple categories, the percentages do not sum to 100%.
+Unfortunately, there is currently no automated tool that can further reliably explain performance speedups at the assembly level.
+Automated analysis and explanation of performance differences in assembly code remains a challenging and unsolved problem.
+The observed gains and our analysis suggest that LLMs can produce reasonable code transformations that meaningfully improve assembly performance.
+These results indicate that LLMs can uncover nuanced optimization opportunities beyond the reach of traditional compiler heuristics, suggesting that compilers themselves still have room for improvement.
+5.4 Comparison with Simple Random Perturbations
+To assess whether the improvements achieved by LLMs stem from systematic strategies learned during training rather than simple random variation in assembly programs, we compare against a baseline that applies simple random perturbations to assembly code.
+While such perturbations do not necessarily improve performance, running them multiple times and selecting the fastest version allows us to compute a speedup.
+These perturbations preserve correctness but can still influence performance, and we report speedup using the minimum observed runtime across 10 runs.
+To randomly perturb the assembly, we compile the binaries with options that randomize both the stack and program load addresses.
+This alters code layout in memory, which can affect performance but not correctness.
+Speedup is measured as the ratio of the fastest run to the original runtime over 10 randomized runs.
+Repeating the experiment five times yields a speedup of 1.19 ± 0.01× (95% CI), significantly lower than 1.46 ± 0.12× achieved by SuperCoder.
+This indicates that SuperCoder systematically exploits performance opportunities rather than relying on simple random perturbations, particularly since it produces each result in a single shot without retries or iterative refinement.
+6 Discussion
+Alternative Reward Function Design.
+Besides the reward function presented in Section˜3.3, we also evaluate an alternative design.
+The original design assigns zero reward whenever any test fails.
+In contrast, the alternative assigns (i) a reward of 1-1 if the program fails to compile, (ii) a partial reward equal to the fraction of passed tests if only some tests succeed, and (iii) a reward of $1+\alpha\cdot\text{speedup}$ once all tests pass.
+Training the base model with this design yields an average speedup of 1.38×.
+Varying the scaling factor (5 or 10) has a negligible effect, and the result remains worse than the 1.46× achieved by SuperCoder with the original reward.
+This suggests that directly optimizing for the terminal speedup reward is more effective, since the alternative design allows models to avoid penalties by merely compiling or even gain positive reward from partially passing tests, which does not translate into actual speedup.
+Direct Compilation from C.
+We also examine a more challenging setting where gcc -O0 assembly is not provided.
+Instead, LLMs receive only the C source code and are asked to generate assembly directly.
+This setup leads to a sharp drop in performance: for example, SuperCoder, which attains 95.0% correctness with the assembly baseline, fails to produce any compilable code without it.
+Similar degradation occurs for other models such as llm-compiler-13b and Claude models.
+These results indicate that, at least in their current state, LLMs can act as superoptimizers building on compiler outputs, but cannot replace compilers themselves.
+Case Study.
+We illustrate in Appendix A.1 a representative example where an LLM discovers an optimization beyond the reach of a state-of-the-art compiler.
 The original C function computes the population count (i.e., the number of set bits) by repeatedly shifting the input and accumulating its least significant bit.
-The assembly code produced by gcc -O3 preserves this loop structure, relying on explicit bitwise operations and conditional branches to compute the result.
-In contrast, the LLM generates a significantly more concise and efficient implementation that replaces the entire loop with a single popcnt instruction.
-This instruction, supported by modern x86-64 architectures, performs the same computation in one operation, thereby reducing both instruction count and runtime overhead.
-Such a transformation is beyond the reach of gcc -O3, which applies a predetermined sequence of rule-based optimization passes and does not conduct semantic-level rewrites of this kind.
-In this case, the language model is able to synthesize functionally equivalent code that exploits hardware-level instructions not utilized by the compiler.
-This demonstrates the potential of language models to optimize assembly by exploring a broader space of semantics-preserving program transformations.
+The assembly produced by gcc -O3 preserves this loop structure, relying on explicit bitwise operations and conditional branches.
+In contrast, Claude-opus-4 produces an efficient implementation that replaces the entire loop with a single popcnt instruction, performing the same computation in one operation and reducing both instruction count and runtime overhead.
+Limitation and Potential Directions.
+A limitation of our approach is the lack of formal correctness guarantees.
+Although we validate generated programs with input–output test cases, testing is inherently incomplete and may miss corner cases even with high coverage.
+However, to our knowledge, no existing verifier can generally handle our assembly programs with loops.
+We hope this work motivates future research in the formal methods communities toward scalable verification techniques.
+The dataset we introduce could serve as a meaningful benchmark for such research.
+Compared with traditional compiler optimization, reinforcement learning introduces substantial computational overhead during training.
+This cost, however, is incurred only once: once trained, the model can generate optimized assembly efficiently with low latency.
+One promising direction for future work is to incorporate an interactive refinement loop, where the model iteratively improves its output based on feedback from errors or performance measurements.
+In addition, most superoptimization research has centered on x86-64, and our work follows this focus.
+Extending the methodology to other targets such as ARM, RISC-V, or GPU kernels could greatly broaden its applicability and impact.
+Such extensions would mainly require adapting the data collection and testing infrastructure, while keeping the core methodology unchanged.
+7 Conclusion
+We investigated whether LLMs can act as superoptimizers, generating assembly programs that outperform code already optimized by industry-standard compilers.
+To support this study, we introduced the first large-scale benchmark of 8,072 assembly programs.
+Evaluating 23 models revealed the difficulty of the task, with most failing to achieve meaningful gains.
+By fine-tuning with reinforcement learning, our model SuperCoder improved both correctness and performance, reaching 95.0% test pass rate and an average speedup of 1.46× over gcc -O3.
+These results demonstrate, for the first time, that LLMs can be applied as superoptimizers for assembly code, opening new opportunities for performance optimization beyond compiler heuristics.
+LLM Usage
+Large Language Models (LLMs) are the primary subject of study in this paper. In addition, we used LLMs as a general-purpose writing assistant to polish the presentation and improve readability.
+Appendix A Appendix
+A.1 Case Study
 C Code
@@
 retq
-LLM Generated
+Claude-Opus-4
 .L0:
@@
 retq
-Figure 2: Case study comparing the C code, baseline assembly produced by gcc -O3, and optimized assembly generated by the LLM.
+Figure A1: Case study comparing the C code, baseline assembly produced by gcc -O3, and optimized assembly generated by Claude-Opus-4.
 The model successfully replaces the loop with the specialized hardware instruction popcnt, resulting in a significantly more concise implementation.
-6 Discussion
-Limitations.
-A key limitation of our approach is the absence of formal correctness guarantees.
-Although we validate generated programs using input-output test cases, such testing is inherently incomplete and may overlook edge cases.
-Consequently, unlike fully verified systems such as Stoke [11], LLM-generated assembly may produce wrong results or exhibit undefined behavior.
-This limitation reflects a broader challenge in programming languages: verifying the semantic equivalence between two arbitrary programs is undecidable in the general case.
-Another limitation lies in the inherent randomness of performance measurement on real hardware.
-Although we mitigate noise through repeated measurements, low-level hardware fluctuations can still introduce variability into speedup estimates.
-Such nondeterminism is difficult to eliminate entirely in real-world settings.
-While prior work has adopted simulator-based evaluation [6], simulators may fail to faithfully capture the actual hardware performance.
-Finally, the observed performance gains may not generalize across machines.
-The model may implicitly learn and exploit hardware-specific characteristics like cache size.
-As a result, a model trained on one machine may not retain its effectiveness when deployed on a different machine.
-Future Work.
-While we use Proximal Policy Optimization (PPO), future work may explore alternative reinforcement learning algorithms such as GRPO [71].
-Expanding to larger and more diverse datasets, particularly those involving performance-critical code beyond competitive programming, would make the setting more realistic and applicable.
-Combining reinforcement learning with supervised fine-tuning may also be beneficial, although it remains unclear whether training on gcc -O3 outputs would provide additional gains.
-Another direction is to incorporate an interactive refinement loop, where the model iteratively updates its output using feedback from errors or performance measurements.
-Finally, extending our approach from x86-64 to other architectures such as MIPS, ARM, or GPU programming could broaden its applicability and impact.
-7 Conclusion
-We explore the use of large language models (LLMs) for optimizing assembly code, a setting where fine-grained control over execution enables performance improvements that are difficult to express in high-level languages.
-While traditional compilers rely on fixed rule-based transformations, they face fundamental limitations due to the complexity of the optimization space.
-To address this, we apply reinforcement learning to fine-tune LLMs with Proximal Policy Optimization (PPO), using a reward function based on correctness and speedup over the gcc -O3 baseline.
-To support this effort, we introduce a benchmark of 8,072 real-world C programs with compiler-generated baseline assembly and test cases.
-Our resulting model, Qwen2.5-Coder-7B-PPO, achieves the highest compile and test pass rates (96.0%) and the best average speedup (1.47×), outperforming all 20 other models evaluated across all metrics.
-These results indicate that reinforcement learning can unlock the potential of LLMs to serve as effective optimizers for assembly code performance.
-Appendix A Appendix
-A.1 Training Configurations
+A.2 Training Configurations
 Component Setting
@@
 Hardware 4× A100 GPUs
-Table A1: Key training configurations for VERL PPO fine-tuning.
-A.2 Prompt Template
+Table A1: Key training configurations for PPO fine-tuning.
+A.3 Prompt Template
 Prompt Template Given the following C code and assembly code, your task is to generate highly optimized x86-64 assembly code.

### v3

@@
 To further enhance performance, we fine-tune models with reinforcement learning, optimizing a reward function that integrates correctness and performance speedup.
-Starting from Qwen2.5-Coder-7B-Instruct (61.4% correctness, 1.10× speedup), the fine-tuned model SuperCoder attains 95.0% correctness and 1.46× average speedup.
+Starting from Qwen2.5-Coder-7B-Instruct (61.4% correctness, 1.10× speedup), the fine-tuned model SuperCoder attains 95.0% correctness and 1.46× average speedup, with additional improvement enabled by Best-of-N sampling and iterative refinement.
 Our results demonstrate, for the first time, that LLMs can be applied as superoptimizers for assembly programs, establishing a foundation for future research in program performance optimization beyond compiler heuristics.
+Machine Learning, ICML
 1 Introduction
@@
 Decades of research have tackled the problem of code optimization, giving rise to two main approaches.
-The first develops better algorithms for rule-based transformations in compilers (Wolf & Lam, 1991).
+The first develops better algorithms for rule-based transformations in compilers (Wolf and Lam, 1991).
 However, given the vast space of possible transformations, compiler-optimized code is not guaranteed to be optimal and often leaves performance untapped (Center et al., 1971; Theodoridis et al., 2022).
@@
 Superoptimization is an aggressive form of program optimization that can outperform compiler-optimized code, yet existing literature has focused on very short, straight-line assembly programs without loops.
-Prior work has primarily relied on CPU-based search heuristics, which fail to scale to larger programs (Schkufza et al., 2013; Phothilimthana et al., 2016; Koenig et al., 2021b); available datasets include at most 15 lines of straight-line assembly (Koenig et al., 2021a).
+Prior work has primarily relied on CPU-based search heuristics, which fail to scale to larger programs (Schkufza et al., 2013; Phothilimthana et al., 2016; Koenig et al., 2021a); available datasets include at most 15 lines of straight-line assembly (Koenig et al., 2021b).
 Figure 1: Overview of the assembly code optimization task.
@@
 Each instance includes input–output test cases and baseline assembly generated by the compiler at its highest optimization level (gcc -O3), which serves as the starting point for further optimization.
-In contrast, the datasets commonly used in the superoptimization community (Warren, 2013; Schkufza et al., 2013; Koenig et al., 2021a) are either extremely limited in size, containing only 25 programs, or consist of toy examples with 2 to 15 instructions, 74% of which have between 3 and 6 instructions and contain no loops.
+In contrast, the datasets commonly used in the superoptimization community (Warren, 2013; Schkufza et al., 2013; Koenig et al., 2021b) are either extremely limited in size, containing only 25 programs, or consist of toy examples with 2 to 15 instructions without loops.
 Our dataset is substantially larger, with assembly programs averaging 130 lines and including loops.
@@
 We use widely adopted algorithms, including Proximal Policy Optimization (PPO) and Group Relative Policy Optimization (GRPO), to train an LLM with a reward function that integrates both correctness and performance speedup.
-Prior work on LLM-based performance optimization has explored alternative methodologies such as supervised fine-tuning (Shypula et al., 2023), chain-of-thought prompting (Liu et al., 2024c), agent-based frameworks (Wei et al., 2025b; c), and preference learning (Du et al., 2024).
-Our approach optimizes speedup explicitly in the reward function, making reinforcement learning well suited to superoptimization.
-To our knowledge, this is the first application of reward-based RL to LLMs for code performance optimization, with correctness and speedup jointly encoded in the objective.
+Prior work on LLM-based performance optimization has explored alternative methodologies such as supervised fine-tuning (Shypula et al., 2024), chain-of-thought prompting (Liu et al., 2024c), agent-based frameworks (Wei et al., 2025b, c), and preference learning (Du et al., 2024).
+Our approach optimizes speedup explicitly in the reward function, making reinforcement learning well-suited to superoptimization.
+To our knowledge, this is among the first applications of reward-based RL to LLMs for code performance optimization, with correctness and speedup jointly encoded in the objective.
 We evaluate 23 LLMs on this task and find that the best-performing model, Claude-opus-4, achieves a 51.5% test-passing rate and an average speedup of 1.43× over the compiler-optimized baseline (gcc -O3).
-Our reinforcement learning approach is highly effective: starting from the base model Qwen2.5-Coder-7B-Instruct, which achieves 61.4% correctness and a modest 1.10× speedup, the fine-tuned model SuperCoder attains 95.0% correctness and 1.46× average speedup.
-We further analyze the code transformation patterns learned by the model.
+Our reinforcement learning approach is highly effective: starting from the base model Qwen2.5-Coder-7B-Instruct, which achieves 61.4% correctness and a modest 1.10× speedup, the fine-tuned model SuperCoder attains 95.0% correctness and 1.46× average speedup, with further improvement enabled by Best-of-N sampling and iterative refinement.
 In summary, our contributions are as follows:
@@
 This far surpasses prior loop-free datasets under 15 lines and marks a substantial step forward in scale and realism for evaluating superoptimization techniques.
-  • We evaluate 23 LLMs on the benchmark and show that RL-based training substantially improves performance: fine-tuning Qwen2.5-Coder-7B-Instruct (61.4% correctness, 1.10× speedup) results in SuperCoder with 95.0% correctness and 1.46× speedup.
+  • We evaluate 23 LLMs on the benchmark and show that RL-based training substantially improves performance: fine-tuning Qwen2.5-Coder-7B-Instruct (61.4% correctness, 1.10× speedup) results in SuperCoder with 95.0% correctness and 1.46× speedup, with further gains enabled by Best-of-N sampling and iterative refinement.
 2 Related Work
@@
 In parallel, many models have been developed to enhance code generation capabilities, such as Codex (Chen et al., 2021), AlphaCode (Li et al., 2022), CodeGen (Nijkamp et al., 2022), InCoder (Fried et al., 2022), StarCoder (Li et al., 2023), DeepSeek-Coder (Guo et al., 2024), Code Llama (Roziere et al., 2023), and others (Hui et al., 2024; Wei et al., 2025e).
-Beyond code generation, LLMs have been applied to real-world software engineering tasks including automated program repair (Xia & Zhang, 2022; Xia et al., 2023b), software testing (Xia et al., 2023a; Deng et al., 2024), bug localization (Yang et al., 2024a), transpilation (Yang et al., 2024c; Bhatia et al., 2024), equivalence checking (Wei et al., 2025a), and synthesis (Wei et al., 2025d).
+Beyond code generation, LLMs have been applied to real-world software engineering tasks including automated program repair (Xia and Zhang, 2022; Xia et al., 2023b), software testing (Xia et al., 2023a; Deng et al., 2024), bug localization (Yang et al., 2024a), transpilation (Yang et al., 2024c; Bhatia et al., 2024), equivalence checking (Wei et al., 2025a), and synthesis (Wei et al., 2025d).
 SWE-bench (Jimenez et al., 2023) integrates these tasks into a benchmark for resolving real GitHub issues.
@@
 CodeRosetta (TehraniJamsaz et al., 2024) targets automatic parallelization, such as translating C++ to CUDA.
-Other efforts focus on optimizing Python code for efficiency (Du et al., 2024; Liu et al., 2024c) or enabling self-adaptation (Huang et al., 2024), and improving C++ performance (Shypula et al., 2023).
+Other efforts focus on optimizing Python code for efficiency (Du et al., 2024; Liu et al., 2024c) or enabling self-adaptation (Huang et al., 2024), and improving C++ performance (Shypula et al., 2024).
 Of particular relevance are approaches to low-level code optimization (Wei et al., 2024; Ouyang et al., 2025).
-The LLM Compiler foundation models (Cummins et al., 2024; 2025) are primarily designed for code size reduction and binary disassembly, whereas our work focuses on optimizing assembly code for performance.
+The LLM Compiler foundation models (Cummins et al., 2024, 2025) are primarily designed for code size reduction and binary disassembly, whereas our work focuses on optimizing assembly code for performance.
 LLM-Vectorizer (Taneja et al., 2025) offers a formally verified solution for auto-vectorization, a specific compiler pass.
@@
 AutoTVM (Chen et al., 2018) pioneered statistical cost model-based search for CUDA code optimization, followed by methods such as Ansor (Zheng et al., 2020), AMOS (Zheng et al., 2022), and others (Shao et al., 2022; Zhao et al., 2024; Wu et al., 2024).
-More recently, LLMs have been explored as code optimizers (Shypula et al., 2023; Grubisic et al., 2024; Wei et al., 2024; 2025b; 2025c), with growing interest in reinforcement learning that guides generation through reward signals (Dou et al., 2024; Wei et al., 2025f).
+More recently, LLMs have been explored as code optimizers (Shypula et al., 2024; Grubisic et al., 2024; Wei et al., 2024, 2025b, 2025c), with growing interest in reinforcement learning that guides generation through reward signals (Dou et al., 2024; Wei et al., 2025f).
 Rewards are often derived from unit-test correctness (Le et al., 2022; Shojaee et al., 2023; Liu et al., 2023a) or binary preference signals (Liu et al., 2024b; Du et al., 2024).
-To our knowledge, this is the first work to apply reinforcement learning to optimize code performance with LLMs, with concurrent efforts exploring CUDA kernel optimization (Li et al., 2025; Baronio et al., 2025).
+To our knowledge, this is among the first works to apply reinforcement learning to optimize code performance with LLMs, with concurrent efforts exploring CUDA kernel optimization (Li et al., 2025; Baronio et al., 2025).
 3 Methodology
@@
 Instead, the model produces arbitrary assembly code, and correctness is validated post hoc via compilation and test execution.
-We train an LLM using reinforcement learning (see Section˜3.3) to generate candidates that both satisfy correctness and achieve performance improvements.
+We train an LLM using reinforcement learning (see Section 3.3) to generate candidates that both satisfy correctness and achieve performance improvements.
 3.2 Dataset Construction
 We construct our dataset using C programs from CodeNet (Puri et al., 2021), a large-scale corpus of competitive programming submissions.
+CodeNet is a well-established and widely used benchmark in the AI-for-code community (Li et al., 2022; Shypula et al., 2024).
 Each dataset instance is a tuple $(C,P,\mathcal{T})$, where $C$ is the original C source code, $P=gcc(C)$ is the corresponding x86-64 assembly generated by compiling $C$ with gcc at the -O3 optimization level, and $\mathcal{T}=\{(x_{i},y_{i})\}_{i=1}^{n}$ is the test set.
@@
 Such a strategy serves two purposes: (1) it favors programs with complex logic that lead to suboptimal performance under -O0 and can be effectively optimized by -O3, and (2) it creates a more challenging setting by starting from code that has already benefited from aggressive compiler optimizations.
-If an LLM can generate code that further improves upon gcc -O3, it suggests that the model can outperform the compiler’s “expert” solution.
-The final dataset consists of 7,872 training programs and 200 held-out evaluation programs, with additional statistics provided in Section˜4.
+The final dataset consists of 7,872 training programs and 200 held-out evaluation programs, with additional statistics provided in Section 4.
+We also evaluate our method on an unbiased, randomly sampled dataset distribution in Appendix A.7.
 3.3 Reinforcement Learning
@@
 We define two auxiliary metrics for computing reward:
-$t(P~),\operatorname{pass}(s,a)=\tfrac{1}{|\mathcal{T}|}\sum_{(x,y)\in\mathcal{T}}\mathbf{1}[\tilde{P}(x)=y],\quad\operatorname{speedup}(s,a)$$t(\tilde{P})$,
+$\operatorname{pass}(s,a)=\tfrac{1}{|\mathcal{T}|}\sum_{(x,y)\in\mathcal{T}}\mathbf{1}[\tilde{P}(x)=y],$
+$t(P~),\operatorname{speedup}(s,a)$$t(\tilde{P})$,
 which denote the fraction of test cases passed and the speedup of the generated program $\tilde{P}$ relative to the baseline $P$. We use the following reward function during training:
@@
 Only when the code compiles and passes all tests is a positive reward assigned, equal to the achieved speedup.
+3.4 Best-of-N Sampling, Supervised Fine-Tuning, and Iterative Refinement
+Best-of-N Sampling.
+Generating multiple candidate programs and selecting the strongest one is a well-established strategy for improving code generation quality (Li et al., 2022; Ehrlich et al., 2025).
+In our setting, the best candidate refers to the program that is correct while achieving the fastest execution time.
+Best-of-N sampling is an inference-time technique that can boost performance, but it incurs additional cost because each candidate is tested at runtime.
+Supervised Fine-Tuning.
+To obtain training targets for supervised fine-tuning, we require reference solutions to the superoptimization task.
+However, superoptimization is inherently open-ended: beyond the compiler baseline, there is no unique ground-truth program, and multiple distinct solutions may exist.
+We apply best-of-8 sampling with the base model and treat the highest-quality candidate for each instance as the ground truth.
+We then fine-tune the model using LoRA (Hu et al., 2022).
+Iterative Refinement.
+Iterative refinement is a complementary inference-time technique that can be applied to any model to further improve its outputs.
+After each trial, we feed back the model’s previous attempt: if the generated program fails to compile or fails any test cases, we include the corresponding compiler errors or test failures in the next prompt; if the model produces a correct program, we also include that successful attempt as part of the prompt.
 4 Experimental Setup
 Dataset.
-We describe our dataset construction approach in Section˜3.2.
-Each instance consists of a C source program $C$, the corresponding gcc -O3 compiled assembly $P$, and a set of test cases $\mathcal{T}$ for correctness evaluation.
-The final dataset contains 7,872 training programs and 200 evaluation programs, with average program lengths and test case counts summarized in Table˜1, and additional analysis below.
-Test Coverage.
-Our dataset includes test cases for every program.
-Rather than relying directly on the original submissions, we re-run each program on its inputs to generate correct outputs, thereby fixing errors in prior datasets.
-The resulting test suites of our evaluation dataset achieve an average of 96.2% line coverage and 87.3% branch coverage, demonstrating high test quality.
 Split # Prog. Avg. Tests Avg. LOC
@@
 Table 1: Dataset statistics across training and evaluation splits. LOC = lines of code.
+We describe our dataset construction approach in Section 3.2.
+Each instance consists of a C source program $C$, the corresponding gcc -O3 compiled assembly $P$, and a set of test cases $\mathcal{T}$ for correctness evaluation.
+The final dataset contains 7,872 training programs and 200 evaluation programs, with average program lengths and test case counts summarized in LABEL:tab:dataset, and additional analysis below.
+Test Coverage.
+Our dataset includes test cases for every program.
+Rather than relying directly on the original submissions, we re-run each program on its inputs to generate correct outputs, thereby fixing errors in prior datasets.
+The resulting test suites of our evaluation dataset achieve an average of 96.2% line coverage and 87.3% branch coverage, demonstrating high test quality.
 Speedup by Compilers.
@@
 Test cases are withheld from the model.
-The model is instructed to generate the optimized assembly code.
 We show the prompt template in Appendix A.3.
@@
 For performance, we measure the relative speedup over the gcc -O3 baseline.
-As defined in Section˜3.1, we assign a default speedup of 1× to any candidate that fails to compile, fails any test case, or is slower than the baseline.
+As defined in Section 3.1, we assign a default speedup of 1× to any candidate that fails to compile, fails any test case, or is slower than the baseline.
 This reflects the practical setting where a system can fall back to the gcc -O3 output, resulting in no performance gain.
@@
 In addition, we include models distilled from DeepSeek-R1 (Guo et al., 2025) based on Qwen and Llama.
-Finally, we evaluate recent compiler-oriented foundation models (Cummins et al., 2024; 2025), pretrained on assembly code and derived from Code Llama, with a design focus on compiler tasks (listed as llm-compiler in Table˜2).
+Finally, we evaluate recent compiler-oriented foundation models (Cummins et al., 2024, 2025), pretrained on assembly code and derived from Code Llama, with a design focus on compiler tasks (listed as llm-compiler in Table 2).
 All open-source models are instruction-tuned.
 Performance Measurement.
-To ensure an accurate performance evaluation, we use hyperfine (hyp, 2025), a benchmarking tool that reduces measurement noise by performing warmup runs followed by repeated timed executions.
+To ensure an accurate performance evaluation, we use hyperfine (27), a benchmarking tool that reduces measurement noise by performing warmup runs followed by repeated timed executions.
 For each program’s execution, we discard the first three runs and report the average runtime over the next ten runs.
@@
 We implement our customized reinforcement learning reward functions within the verl framework (Sheng et al., 2024), which enables fine-tuning of LLMs using PPO and GRPO.
-As part of this setup, we build a task-specific environment that handles program compilation, test execution, and runtime measurement, as detailed in Section˜3.3.
+As part of this setup, we build a task-specific environment that handles program compilation, test execution, and runtime measurement, as detailed in Section 3.3.
 This environment provides the model with direct scalar feedback based on both functional correctness and execution performance.
 Training Configurations.
-Among all evaluated models (see Table˜2), we select Qwen2.5-Coder-7B-Instruct for training due to its strong correctness results and substantial room for performance improvement, while intentionally avoiding compiler-specific foundation models to preserve generality.
+Among all evaluated models (see Table 2), we select Qwen2.5-Coder-7B-Instruct for training due to its strong correctness results and substantial room for performance improvement, while intentionally avoiding compiler-specific foundation models to preserve generality.
 Training is performed on a single node with four A100 GPUs.
-Full hyperparameter settings are provided in Appendix A1.
+Hyperparameter settings are in Appendix A.2.
 5 Results
@@
 We report compilation success rate, test pass rate, and average speedup over the gcc -O3 baseline.
-All open-source models are instruction-tuned.
+Speedup is averaged across all test inputs, with each input evaluated over ten runs.
 Main Results.
-Table˜2 reports results across evaluated models.
+Table 2 reports results across evaluated models.
 Most perform poorly on this task, with only a few demonstrating effectiveness as superoptimizers.
@@
 Improvement.
-Table˜3 presents the results of RL training, averaged over 5 runs with 95% confidence intervals to provide more statistical confidence in the reported improvements.
+Table 3 presents the results of RL training, averaged over 5 runs with 95% confidence intervals to provide more statistical confidence in the reported improvements.
 We select Qwen2.5-Coder-7B-Instruct for RL training due to its strong test pass rate (61.4%) among models.
@@
 […]
-Table 3: Performance of the base model and the models fine-tuned with RL. Results include compilation success, test pass rates, and average speedup, reported over 5 runs with 95% confidence intervals.
-5.3 Analysis of Learned Program Transformations
-To better understand why LLMs can further optimize assembly programs already optimized by industry-standard compilers, we analyze all 200 evaluation programs by comparing the gcc -O3 output with the assembly generated by our PPO-trained model SuperCoder.
-We categorize the learned transformations into four types: (1) code layout and instruction scheduling, which account for 98.5% of the changes and include basic block placement, instruction alignment, and reordering to improve cache behavior and hide latencies; (2) register allocation, representing 16% of the changes, where the model selects alternative registers; (3) control flow optimization (1.5%); and (4) instruction selection (1.0%).
-Since a single problem’s transformation may belong to multiple categories, the percentages do not sum to 100%.
-Unfortunately, there is currently no automated tool that can further reliably explain performance speedups at the assembly level.
-Automated analysis and explanation of performance differences in assembly code remains a challenging and unsolved problem.
-The observed gains and our analysis suggest that LLMs can produce reasonable code transformations that meaningfully improve assembly performance.
-These results indicate that LLMs can uncover nuanced optimization opportunities beyond the reach of traditional compiler heuristics, suggesting that compilers themselves still have room for improvement.
-5.4 Comparison with Simple Random Perturbations
-To assess whether the improvements achieved by LLMs stem from systematic strategies learned during training rather than simple random variation in assembly programs, we compare against a baseline that applies simple random perturbations to assembly code.
-While such perturbations do not necessarily improve performance, running them multiple times and selecting the fastest version allows us to compute a speedup.
-These perturbations preserve correctness but can still influence performance, and we report speedup using the minimum observed runtime across 10 runs.
-To randomly perturb the assembly, we compile the binaries with options that randomize both the stack and program load addresses.
-This alters code layout in memory, which can affect performance but not correctness.
-Speedup is measured as the ratio of the fastest run to the original runtime over 10 randomized runs.
-Repeating the experiment five times yields a speedup of 1.19 ± 0.01× (95% CI), significantly lower than 1.46 ± 0.12× achieved by SuperCoder.
-This indicates that SuperCoder systematically exploits performance opportunities rather than relying on simple random perturbations, particularly since it produces each result in a single shot without retries or iterative refinement.
+[…]
+Table 3: Performance of the base model and the models trained with RL or supervised fine-tuning.
+Results include compilation success, test pass rates, and average speedup, reported over 5 runs with 95% confidence intervals.
+5.3 Results from Supervised Fine-Tuning and Inference-Time Methods
+Best-of-N Sampling.
+We evaluate best-of-N sampling for three models: claude-opus-4 (the strongest baseline in Table 2), Qwen2.5-Coder-7B (base), and SuperCoder (our PPO-trained model).
+Results are shown in Figure 3.
+Notably, the base model’s best-of-8 speedup is close to the PPO-trained model’s best-of-1 result, and the RL-trained model itself can still be improved with best-of-N sampling (i.e., from 1.46× in the single-sample setting to 1.93× with best-of-8 sampling).
+Supervised Fine-Tuning.
+We describe our supervised fine-tuning approach in Section 3.4.
+Table 2 reports results averaged over five runs with 95% confidence intervals.
+While supervised fine-tuning improves performance, RL achieves slightly stronger results.
+We believe that RL is a natural fit for the open-ended nature of superoptimization, as RL directly optimizes for correctness and speedup rather than imitating existing examples.
+Iterative Refinement.
+Figure 3 shows the results of iterative refinement, where the model receives feedback about compilation failures, test failures, or performance for self-reflection.
+All three models exhibit improvements as the number of refinement iterations increases, with the effect being most pronounced for our RL fine-tuned model.
+Figure 2: Best-of-N sampling results.
+Figure 3: Iterative refinement results.
+5.4 Analysis of Program Transformations
+To better understand why LLMs can further optimize assembly programs already optimized by industry-standard compilers, we analyze all 200 evaluation programs by comparing the gcc -O3 output with the assembly generated by claude-opus-4 with best-of-8 sampling.
+We categorize the program transformations into nine types: 1) Loop restructuring covers changes to loop organization, including reordering, unrolling, or altered control flow; 2) Instruction selection captures the use of specialized CPU instructions (e.g., bsrq, popcnt, cmov) in place of longer generic instruction sequences; 3) Algorithmic simplification denotes replacing complex or custom logic with simpler algorithms or standard library routines such as memcmp, strcmp, or atoi; 4) Stack canary removal refers to eliminating stack protection checks and related security instrumentation; 5) Register allocation reflects differences in how registers are assigned, reused, or spilled to memory; 6) Branch elimination involves removing conditional branches by using conditional moves (cmov) or condition-setting instructions (setcc); 7) Address calculation optimization refers to simplifying memory address computations and offset arithmetic; 8) Function call optimization captures substituting heavyweight or checked function variants with simpler equivalents (e.g., __isoc99_scanf vs.
+scanf); and 9) Arithmetic optimization includes simplifying arithmetic operations such as replacing divisions or multiplications with shifts, or using increment/decrement instructions where applicable.
+Appendix A.6 presents a detailed category-wise analysis, reporting the frequency of each transformation category among all transformed programs that pass all tests.
+For a given optimized version, the model may induce transformations spanning multiple optimization categories.
+Because performance optimization typically involves multiple interrelated decisions, it is difficult to automatically isolate which types of program transformations are responsible for observed performance improvements.
+More broadly, automated analysis and attribution of performance differences at the assembly level remains an open problem.
 6 Discussion
+Prompt Optimization Methods.
+We experimented with few-shot in-context learning and found that adding more examples does not reliably improve performance and often degrades it, consistent with prior observations in code optimization (Shypula et al., 2024).
+We also evaluated GEPA (Agrawal et al., 2025), an evolutionary prompting framework that uses natural language reflection to derive optimization rules, but observed only minimal gains.
+Additional results are provided in Appendix A.4.
+Besides, we also experimented with prompts that include only the C source file, without providing the compiler-generated assembly.
+We find that omitting the assembly as a starting point leads to substantially worse results; these findings are discussed in Appendix A.5.
 Alternative Reward Function Design.
-Besides the reward function presented in Section˜3.3, we also evaluate an alternative design.
+Besides the reward function presented in Section 3.3, we also evaluate an alternative design.
 The original design assigns zero reward whenever any test fails.
@@
 Varying the scaling factor (5 or 10) has a negligible effect, and the result remains worse than the 1.46× achieved by SuperCoder with the original reward.
-This suggests that directly optimizing for the terminal speedup reward is more effective, since the alternative design allows models to avoid penalties by merely compiling or even gain positive reward from partially passing tests, which does not translate into actual speedup.
-Direct Compilation from C.
-We also examine a more challenging setting where gcc -O0 assembly is not provided.
-Instead, LLMs receive only the C source code and are asked to generate assembly directly.
-This setup leads to a sharp drop in performance: for example, SuperCoder, which attains 95.0% correctness with the assembly baseline, fails to produce any compilable code without it.
-Similar degradation occurs for other models such as llm-compiler-13b and Claude models.
-These results indicate that, at least in their current state, LLMs can act as superoptimizers building on compiler outputs, but cannot replace compilers themselves.
-Case Study.
-We illustrate in Appendix A.1 a representative example where an LLM discovers an optimization beyond the reach of a state-of-the-art compiler.
-The original C function computes the population count (i.e., the number of set bits) by repeatedly shifting the input and accumulating its least significant bit.
-The assembly produced by gcc -O3 preserves this loop structure, relying on explicit bitwise operations and conditional branches.
-In contrast, Claude-opus-4 produces an efficient implementation that replaces the entire loop with a single popcnt instruction, performing the same computation in one operation and reducing both instruction count and runtime overhead.
+This suggests that directly optimizing for the terminal speedup reward is more effective.
+Case Studies.
+We illustrate in Appendix A.1 two representative examples where an LLM discovers optimizations beyond the reach of a state-of-the-art compiler.
+The first example is about instruction selection.
+In the first example, the original C function computes the population count (i.e., the number of set bits) by repeatedly shifting the input and accumulating its least significant bit.
+The assembly produced by gcc -O3 preserves this loop structure.
+In contrast, Claude-opus-4 produces an efficient implementation that replaces the entire loop with a single popcnt instruction, performing the same computation in one operation.
+The second example involves loop restructuring, along with function call optimization by replacing checked function variants with simpler equivalents.
 Limitation and Potential Directions.
-A limitation of our approach is the lack of formal correctness guarantees.
-Although we validate generated programs with input–output test cases, testing is inherently incomplete and may miss corner cases even with high coverage.
-However, to our knowledge, no existing verifier can generally handle our assembly programs with loops.
-We hope this work motivates future research in the formal methods communities toward scalable verification techniques.
-The dataset we introduce could serve as a meaningful benchmark for such research.
-Compared with traditional compiler optimization, reinforcement learning introduces substantial computational overhead during training.
-This cost, however, is incurred only once: once trained, the model can generate optimized assembly efficiently with low latency.
-One promising direction for future work is to incorporate an interactive refinement loop, where the model iteratively improves its output based on feedback from errors or performance measurements.
-In addition, most superoptimization research has centered on x86-64, and our work follows this focus.
-Extending the methodology to other targets such as ARM, RISC-V, or GPU kernels could greatly broaden its applicability and impact.
-Such extensions would mainly require adapting the data collection and testing infrastructure, while keeping the core methodology unchanged.
+Our evaluation relies on test-based validation, which is common in prior work (Shypula et al., 2024; Du et al., 2024).
+While effective in practice, this approach would be more faithful with access to a formal equivalence checker for assembly programs.
+Existing state-of-the-art equivalence checkers for assembly (Schkufza et al., 2013) cannot handle programs with the rich control flow present in our dataset.
+To mitigate this limitation, our evaluation achieves 96.2% line coverage, substantially reducing the risk of undetected errors.
+Encouragingly, recent work has begun developing formal verification tools for GPU kernels (Dubey et al., 2025), partly motivated by advances in LLM-generated GPU code.
+We hope that our work further motivates the development of more general and scalable verification tools for assembly programs.
+In addition, most superoptimization research has centered on x86-64.
+Extending the methodology to other targets such as ARM, RISC-V, or GPU kernels could greatly broaden its applicability.
 7 Conclusion
@@
 By fine-tuning with reinforcement learning, our model SuperCoder improved both correctness and performance, reaching 95.0% test pass rate and an average speedup of 1.46× over gcc -O3.
+We also show that Best-of-N sampling and iterative refinement can bring additional improvement.
 These results demonstrate, for the first time, that LLMs can be applied as superoptimizers for assembly code, opening new opportunities for performance optimization beyond compiler heuristics.
-LLM Usage
-Large Language Models (LLMs) are the primary subject of study in this paper. In addition, we used LLMs as a general-purpose writing assistant to polish the presentation and improve readability.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
 Appendix A Appendix
-A.1 Case Study
+A.1 Case Studies
 C Code
@@
 The model successfully replaces the loop with the specialized hardware instruction popcnt, resulting in a significantly more concise implementation.
+GCC -O3 Output
+testl %eax, %eax
+jle .L2
+movl %eax, %r8d
+jmp .L5
+.p2align 4,,10
+.p2align 3
+.L4:
+subl \$1, %r8d
+je .L2
+.L5:
+movl %ecx, %eax
+cltd
+idivl %r8d
+testl %edx, %edx
+jne .L4
+movl %r8d, %edx
+leaq .LC1(%rip), %rsi
+movl \$1, %edi
+xorl %eax, %eax
+call __printf_chk@PLT
+.L2:
+movq 8(%rsp), %rax
+subq %fs:40, %rax
+jne .L14
+xorl %eax, %eax
+addq \$24, %rsp
+.cfi_remember_state
+.cfi_def_cfa_offset 8
+ret
+Claude-Opus-4
+testl %eax, %eax
+jle .L2
+.L3:
+movl %ecx, %edx
+movl %eax, %esi
+movl %edx, %eax
+cltd
+idivl %esi
+testl %edx, %edx
+je .L4
+decl %esi
+movl %esi, %eax
+jg .L3
+.L2:
+xorl %eax, %eax
+addq \$24, %rsp
+ret
+.L4:
+leaq .LC1(%rip), %rdi
+xorl %eax, %eax
+call printf@PLT
+xorl %eax, %eax
+addq \$24, %rsp
+ret
+.size main, .-main
+Figure A2: Case study comparing the baseline assembly code snippet produced by gcc -O3 and the optimized assembly code snippet generated by Claude-Opus-4.
+Claude-Opus-4 eliminates the entry-path unconditional jump and alignment padding by fusing GCC’s two-block loop into a single, simpler control-flow structure, calls printf@PLT directly (a simpler function variant without security checks), and removes gcc’s stack-protector canary check.
 A.2 Training Configurations
@@
 Only output the optimized assembly code. Do not include any other text. Do not write any comments in the assembly code. Wrap the assembly code in assembly tags. Optimized Assembly Code:
+A.4 Prompt Optimization Methods
+Few-shot Prompting.
+We evaluate 0-shot, 2-shot, and 4-shot prompting across three models.
+Table A2 shows that adding more in-context examples does not reliably improve performance and often degrades it.
+This observation is consistent with prior work (Shypula et al., 2024), which finds that few-shot examples can bias the model and lead to incorrect task understanding.
+Model Shots Compile Pass (%) Test Pass (%) Avg. Speedup
+claude-opus-4 0-shot 90.0 51.5 1.43×
+claude-opus-4 2-shot 95.0 15.0 1.13×
+claude-opus-4 4-shot 95.0 12.5 1.11×
+SuperCoder (PPO) 0-shot 96.0 95.0 1.46×
+SuperCoder (PPO) 2-shot 94.0 90.5 1.59×
+SuperCoder (PPO) 4-shot 93.0 81.0 1.54×
+Qwen2.5-Coder-7B (Base) 0-shot 77.9 61.4 1.10×
+Qwen2.5-Coder-7B (Base) 2-shot 70.5 35.0 1.10×
+Qwen2.5-Coder-7B (Base) 4-shot 80.5 30.5 1.06×
+Table A2: Comparison of 0-shot, 2-shot, and 4-shot prompting across different models.
+Figure A3: Categorization of the program transformations: Loop Restructuring (LR), Arithmetic Optimization (AO), Address Calculation (AC), Stack Canary Removal (SCR), Algorithmic Simplification (AS), Register Allocation (RA), Function Call Optimization (FC), Instruction Selection (IS), Branch Elimination (BE).
+Prompt Evolution Framework.
+We experimented with GEPA (Agrawal et al., 2025), an evolutionary prompting framework that uses natural language reflection to learn high-level optimization rules from trial and error.
+We used gpt-4o as the model under evaluation and gpt-5 as the reflection model.
+GEPA yielded only modest gains: compilation pass increased from 81.0% to 84.0% and test pass from 5.0% to 7.5%, while performance speedup remained essentially unchanged.
+We suspect this is because assembly optimization requires substantial domain knowledge that is difficult to capture by modifying the prompt alone.
+A.5 Direct Compilation from C
+We also examine a more challenging setting where gcc -O3 assembly is not provided.
+Instead, LLMs receive only the C source code and are asked to generate assembly directly.
+This setup leads to a sharp drop in performance: for example, SuperCoder, which attains 95.0% correctness with the assembly baseline, fails to produce any compilable code without it.
+Similar degradation occurs for other models such as llm-compiler-13b and Claude models.
+These results indicate that, at least in their current state, LLMs can act as superoptimizers building on compiler outputs, but cannot replace compilers themselves.
+A.6 Categorization of Program Transformations
+We categorize the differences between the code optimized by claude-opus-4 with best-of-8 sampling and the compiler-generated code into nine classes, each capturing a distinct kind of optimization or transformation: 1) Loop Restructuring (LR) covers changes to loop organization, including reordering, unrolling, or altered control flow; 2) Arithmetic Optimization (AO) includes simplifying arithmetic operations such as replacing divisions or multiplications with shifts, or using increment/decrement instructions where applicable; 3) Address Calculation (AC) optimization refers to simplifying memory address computations and offset arithmetic; 4) Stack Canary Removal (SCR) refers to eliminating stack protection checks and related security instrumentation; 5) Algorithmic Simplification (AS) denotes replacing complex or custom logic with simpler algorithms or standard library routines such as memcmp, strcmp, or atoi; 6) Register Allocation (RA) reflects differences in how registers are assigned, reused, or spilled to memory; 7) Function Call (FC) optimization captures substituting heavyweight or checked function variants with simpler equivalents (e.g., __isoc99_scanf vs.
+scanf); 8) Instruction Selection (IS) captures the use of specialized CPU instructions (e.g., bsrq, popcnt, cmov) in place of longer generic instruction sequences; 9) Branch Elimination involves removing conditional branches by using conditional moves (cmov) or condition-setting instructions (setcc).
+In  Figure A3, we plot the percentage of instances in our evaluation dataset that have each of these categories.
+Loop Restructuring (LR) is the most frequent category, appearing in 68.2% of instances, indicating that Claude often performs high-level changes to loop structure and control flow.
+Arithmetic Optimization (AO) and Address Calculation (AC) are also common, occurring in 45.0% and 30.5% of cases, respectively, while Stack Canary Removal (SCR) and Algorithmic Simplification (AS) appear in roughly a quarter of the dataset.
+Note that the total percentage exceeds 100% because an instance may have multiple transformation types.
+A.7 Additional Evaluation on Randomly Sampled Programs
+Our main dataset samples programs with larger speedup from -O0 to -O3 speedup.
+To evaluate how much dataset distribution affects our method, we conduct an additional experiment on 200 programs randomly sampled from CodeNet, and we ensure that this new set is fully disjoint from both our training and evaluation splits and not selected based on compiler speedup characteristics.
+Across these unbiased sampled programs, the conclusions remain largely unchanged: the PPO-trained model consistently produces correct and faster assembly code.
+This indicates that the improvements learned through reinforcement learning generalize beyond the specific distribution used for training and primary evaluation.
+Model Compile Pass (%) Test Pass (%) Avg. Speedup
+[…]
+[…]
+[…]
+Table A3: Performance on 200 randomly sampled programs from CodeNet, disjoint from all training and evaluation splits.
+Unlike the main dataset, these programs are not selected for large gcc -O0 to gcc -O3 speedup.
+The PPO-trained model maintains strong correctness and performance gains.